### v1

# LLMAID: Identifying AI Capabilities in Android Apps with LLMs

Abstract.
Recent advancements in artificial intelligence (AI) and its widespread integration into mobile software applications have received significant attention, highlighting the growing prominence of AI capabilities in modern software systems. However, the inherent hallucination and reliability issues of AI continue to raise persistent concerns. Consequently, application users and regulators increasingly ask critical questions such as: “Does the application incorporate AI capabilities?” and “What specific types of AI functionalities are embedded?” Preliminary efforts have been made to identify AI capabilities in mobile software; however, existing approaches mainly rely on manual inspection and rule-based heuristics. These methods are not only costly and time-consuming but also struggle to adapt advanced AI techniques. Moreover, they lack the granularity needed to accurately characterize the diverse range of AI functionalities. Specifically, existing methods often focus on specific types of information, such as AI packages and on-device models, while overlooking other potential AI factors, i.e., APIs and HTTPS requests. As a result, these methods fail to provide a comprehensive AI service summarization of software applications.
To address the limitations of existing methods, we propose LLMAID (Large Language Model for AI Discovery), a semantics-driven framework that combines program analysis techniques with large language models (LLMs) to identify and analyze AI services in mobile applications. LLMAID includes four main tasks: (1) candidate extraction, (2) knowledge base interaction, (3) AI capability analysis and detection, and (4) AI service summarization. We apply LLMAID to a dataset of 4,201 Android applications and demonstrate that it identifies 242% more real-world AI apps than state-of-the-art rule-based approaches. Our experiments show that LLM4AID achieves high precision and recall, both exceeding 90%, in detecting AI-related components. Additionally, a user study indicates that developers find the AI service summaries generated by LLMAID to be more informative and preferable to the original app descriptions. Finally, we leverage LLMAID to perform an empirical analysis of AI capabilities across Android apps. The results reveal a strong concentration of AI functionality in computer vision (54.80%), with object detection emerging as the most common task (25.19%).
1. Introduction
Advancements in artificial intelligence (AI) and their successful integration into mobile applications (apps) have enhanced the capabilities of Android apps in daily activities, such as optical character recognition, autonomous driving, object recognition and shopping recommendations. To support developers’ integration of AI capabilities, several vendors have introduced mobile AI frameworks, such as TensorFlow Lite (abadi2016tensorflow), Caffe2 (caffe2:2023), Core ML (coreml:2023), and ncnn (ncnn:2023). These frameworks enable efficient on-device model building and inference, offering better privacy protection and reducing bandwidth usage compared to cloud-based models (guo2018cloud), while also improving responsiveness and ensuring functionality in resource-constrained environments typical of mobile devices. In recent years, there has been a surge in the development of large AI models, such as stable diffusion (stable:2023), ChatGPT (chatgpt:2023), LLaMA (touvron2023llama), and DeepSeek (liu2024deepseek). Unlike traditional lightweight frameworks, these models provide more advanced capabilities in reasoning, multimodal content understanding, and generative tasks. With their advanced abilities to understand and generate content, emerging app developers increasingly resort to these models via the cloud to enhance their businesses (xu2019first; hu2023first), enabling the creation of richer user experiences such as personalized recommendations, interactive assistants, and multimodal content generation.
With the widespread adoption of AI and the growing awareness of its inherent limitations, issues of AI discoverability and transparency have attracted attention from both application consumers and various regulatory bodies. To have a comprehensive understanding of AI services in software artifacts for these stakeholders, this study first tries to uncover the current status quo of AI services in real-world Android applications. In this paper, we define this process as AI Discovery, which encompasses determining whether AI capabilities are integrated or not and further elucidating the functional characteristics of the incorporated AI services. Several studies have been conducted to investigate AI techniques in smartphone apps (xu2019first; zhao2022survey; li2022ai; siu2023towards; deng2019deep). Specifically, Deng et al. (deng2019deep) and Zhao et al. (zhao2022survey) both provided comprehensive surveys on deep learning for mobile devices, highlighting hardware architectures, software frameworks, and optimization mechanisms such as pruning, quantization, and hardware accelerators that enable efficient, privacy-preserving, and cost-effective on-device AI across diverse application domains. Siu et al. (siu2023towards) empirically investigated user–AI interaction in 176 mobile apps, identifying 255 AI features and 759 implementations, ultimately categorizing interaction patterns and demonstrating the importance of usable interaction design for on-device AI features. Li et al. (li2022ai) conducted the largest empirical study to date, analyzing 56,682 AI apps with their newly proposed rule-based approach, analyzing application trends, framework adoption, model protection, and user-related concerns, and releasing an open dataset for future research. Xu et al. (xu2019first) examined 16,500 popular Android apps to reveal early adopters of deep learning, their usage scenarios, and model characteristics, offering insights that emphasize both the potential of mobile deep learning and the need for optimizations, model protection, and validation.
However, these studies exhibit several notable constraints: 1) lacking attention to other potential AI components, [^1] An AI component may refer to either high-level modules (e.g., task planning engines) or low-level implementation artifacts (e.g., APIs, libraries, HTTP requests), whereas an AI feature denotes high-level functionalities (e.g., image recognition, sentiment analysis) enabled by these components. such as integrated development packages, internal assets, and HTTP requests and API invocations in detailed implementations and integrated libraries. Ignoring these components would lead to an incomplete and insufficient analysis; 2) relying on manual inspection or pre-defined keyword searches, which require up-to-date human expertise, making them expensive and difficult to scale; 3) depending on human-defined rule-based methods, which are constrained by domain knowledge and thus ineffective for identifying AI components in diverse real-world applications.
To address these aforementioned challenges, we argue that automating the human-like semantic-based AI Discovery is essential, enabling the characterization of real-world AI components by interpreting their meaningful names. Inspired by recent findings (hagendorff2023human; ning2024can; li2024llms; pan2024large) that large language models (LLMs) are capable of exhibiting human-like reasoning and decision-making, we explore how these models can enhance the AI Discovery process by interpreting semantically meaningful AI component names. Moreover, recent studies have shown that these LLMs can automate various software tasks, including automatic program repair (zhang2022repairing; jiang2023impact; xia2023conversational; han2024chase), code generation (nijkamp2022conversational; chakraborty2022natgen; fried2022incoder; liao20243), and improving software analysis efficiency and effectiveness (ma2023scope; li2023hitchhiker). These findings provide evidence that LLMs may have substantial knowledge in software practice, motivating us to explore their full potential in AI Discovery within mobile applications.
To this end, we propose LLMAID (Large Language Model for AI Discovery), the first framework that leverages large language models (LLMs) to identify AI components extracted through static program analysis and interpret their functional capabilities. The LLMAID framework consists of four key stages: (1) candidate extraction, (2) knowledge base interaction, (3) AI capability analysis and detection, and (4) AI service summarization. This structured approach allows us to perform AI discovery in an automated, step-by-step manner, significantly reducing the need for manual intervention.
We implemented LLMAID using the widely adopted GPT-3.5-turbo model, which is much more cost-effective than the GPT-4 model (gptmodel:2023). To evaluate the effectiveness of LLMAID, we conducted validation on a sample of 382 AI applications selected from a total of 56,682 already identified AI apps (li2022ai). Manual evaluation of the results confirms that LLMAID achieves high precision and recall (exceeding 90%) in detecting AI components. Furthermore, we applied LLMAID to a dataset of 3,819 real-world Android apps published on Google Play between 2019 and 2023. Compared to the state-of-the-art (SOTA) detection approach, AI Discriminator (li2022ai), LLMAID identifies more than twice as many AI Android apps, showcasing its superior detection capability. Our analysis reveals that the majority of detected AI components pertain to the domains of computer vision (54.80%) and data analysis (26.85%), with object detection (25.19%) and data processing (23.17%) being the most prevalent tasks. At the application level, we find that 47.44% (643/1366) of AI-enabled apps offer services related to data analysis, while 46.41% (634/1366) incorporate computer vision functionalities. We further demonstrate the practical utility of LLMAID by evidencing that developers perceive its AI service report as both informative and preferable.
Our main contributions are as follows:
  • We present LLMAID, a prototype that automatically identifies AI components in Android apps and generates corresponding AI functionality summarizations.
  • We curate datasets covering both AI apps and fine-grained AI components to support future research.
  • We conduct the first analysis to uncover prevalent AI domains and tasks within Android apps.
Figure 1. The workflow of LLMAID.
2. Approach
Fig. 1 illustrates the general framework and workflow of LLMAID (LLM for AI Discovery). As mentioned above, the LLMAID framework consists of four stages: candidate extraction, knowledge base interaction, AI capability analysis and detection, and AI service summarization. Together, these stages are designed to address the complexity of uncovering AI functionalities in Android apps. In the first stage, Candidate extraction, static analysis techniques are applied to the Android APK to extract fine-grained components that may potentially involve AI functionality. The second stage, Knowledge base interaction, refines these candidates by referencing dynamically constructed AI and non-AI knowledge bases, which act as caches of previously analyzed components. This step reduces redundant analysis and improves precision by filtering out irrelevant components early. The third stage, AI capability analysis and detection, prompts a large language model to interpret the functionality of each remaining candidate and determine whether it constitutes an AI component, yielding a curated set of AI-relevant elements together with functional descriptions. Finally, the AI service summarization stage produces a concise AI service report for the given APK by leveraging a carefully designed LLM prompt, which incorporates the determined AI components along with their interpreted AI functional descriptions. Importantly, the overall workflow is model-agnostic and can be applied with any advanced large language model, ensuring adaptability as LLM technologies continue to evolve. We elaborate on each stage in the following subsections.
Figure 2. The structure of a typical released Android APK.
2.1. Task 1: Candidate Extraction
To determine AI capabilities within an Android application, we first extract all relevant information—referred to as candidates-from the APK file. Fig. 2 shows the structure of a typical released Android APK file, typically including classes.dex, which is the Dalvik executable file (dalvik:2022), assets, META-INFO, lib, AndroidManifest.xml, res, and resources.arsc. Notably, an APK may comprise multiple .dex files because each file is constrained by the 65,536-method limit. When this threshold is exceeded, the compiler automatically partitions the code into additional files, conventionally named classes1.dex, classes2.dex, and so forth.
We utilize AAPT (aapt:2024) to extract the structure from the curated Android APK file. Our focus is on examining the implementation details embedded in the app, which requires decompiling .dex files to retrieve the underlying Java logic and analyzing .so files in the lib directory for native code. Specifically, we extend a static analysis approach, namely App Scanner, proposed by Liu et al. (liu2021identifying) to extract APIs from an APK file. The scanner leverages the Soot framework (lam2011soot), a widely used Java-based static analysis framework that has been extended to support Android applications. In particular, Soot translates Dalvik bytecode (dalvik:2022), the default executable format for Android applications generated from Java (or Kotlin) source code through an intermediate compilation step and optimized for devices with limited memory and power resources, into Jimple (bartel2012dexpler). Jimple is a three-address intermediate representation that is both human-readable and well suited for static program analysis. This transformation bridges the gap between low-level Dalvik instructions and a higher-level abstraction and enables a more effective analysis of Android applications. The App Scanner method in this study is enhanced to additionally capture fine-grained components from multiple sources within the given APK, such as .dex files, native library .so files, and the asset files. This also allows us to extract more indicators of AI functionality, such as HTTP requests https://api.openai.com/v1/ and descriptive package name com.google.firebase.ml.vision.objects. Moreover, to analyze embedded dynamic native libraries, we employ tools readelf (readelf:2023) and objdump (objdump:2023) to parse HTTPS requests, as many apps implement AI capabilities by invoking these requests via the Android NDK (ndk:2023).
2.2. Task 2: Knowledge Base Interaction
Given our focus on implementation-level details of AI capabilities, the extraction process produces a substantial number of fine-grained components that include both AI-relevant and non-AI functionalities. The presence of non-AI components introduces noise into the analysis, as they do not contribute to the understanding of AI capabilities and can impede the accurate identification of AI-related functionality. To reduce such noise, we first apply a rule-based filtering approach designed to exclude non-AI components, such as official Java and Android kernel-related; however, because this method operates at a coarse granularity, some non-AI components may still be incorrectly retained as AI-related.
To address the limitations of coarse-grained rule-based filtering and further improve analysis efficiency, we introduce a knowledge base (KB) interaction step into the workflow. This step enables the reuse of analysis outcomes for components that have already been processed, allowing previously verified AI or non-AI components to be retrieved directly and thereby reducing redundant computation. To support this process, we construct two complementary KBs that function as caches rather than shortcuts for classification: an AI KB, which stores verified AI components together with their capability analyses, and a non-AI KB, which records confirmed non-AI components to facilitate pre-analysis filtering. After component extraction and initial filtering, the system checks each candidate against both KBs; if a match is found, the corresponding analysis and detection result can be directly retrieved, thereby bypassing further inference. Otherwise, the components are forwarded to the next stage, Task 3: AI capability analysis and detection, whose results are then systematically incorporated into the corresponding KB for future reference and reuse. To ensure reliability, both KBs are continuously refined and manually audited, ensuring that cached results maintain high accuracy while significantly reducing unnecessary computation.
2.3. Task 3: AI Capability Analysis and Detection
This stage, the core of LLMAID, focuses on identifying which of the extracted components are AI-related and characterizing them in terms of their AI functional capabilities. To ensure precision and interpretability, this task is divided into two sub-steps: AI capability analysis and AI component determination. The first sub-step, AI capability analysis, interprets the functional behavior of each candidate component to identify the type of functionality it provides, such as image recognition, language processing, or network communication. The second sub-step, AI component determination, then decides whether the interpreted component should be classified as AI-related, filtering out non-AI elements. After the knowledge base interaction phase, we obtain a set of deduplicated candidate components that serve as the input for this two-step process. For each candidate, we issue an initial prompt to ChatGPT to generate a functional-oriented interpretation that supports both capability analysis and component determination. For example, given the package name com.google.mlkit.vision.objects, ChatGPT may provide the analysis result of “Google’s ML Kit API for object detection and tracking in images and videos.” In cases where components are obfuscated or highly customized, i.e., nlp.WordPieceModelPB, ChatGPT can still offer meaningful functional descriptions, e.g., “A library for tokenizing text using the WordPiece algorithm, implemented in Protocol Buffers format” which would be difficult to identify using a rule-based approach. Based on the generated analysis, we further prompt ChatGPT to determine whether the component falls within the scope of AI and justify its decision with a detailed explanation. Leveraging ChatGPT’s ability to interpret and reason over its own generated content, the sub-steps of AI capability analysis and AI component determination can be straightforwardly and effectively conducted by adopting the prompts of Task 3.1 and Task 3.2 in Fig. 3.
2.4. Task 4: AI Service Summarization
Figure 3. Prompt design and examples for LLMAID
Given the determined AI components together with their corresponding functional descriptions, we construct and employ a carefully designed ChatGPT prompt to generate a concise AI service report. This report synthesizes the functional interpretations of the identified components into a coherent summary of their AI capabilities, thereby providing an interpretable overview that bridges low-level implementation details with higher-level functional insights. Nevertheless, due to the inherent context-length limitations of ChatGPT, large language models (LLMs) in general, it is often infeasible to incorporate all component names and their corresponding analyses into a single prompt, which necessitates additional strategies to manage input complexity. To mitigate this constraint, we exclude low-level API components whose functionalities can generally be inferred from their package context, and instead concentrate on higher-level components. The remaining AI-relevant components and their interpreted AI functional descriptions are concatenated and incorporated into a carefully constructed prompt for ChatGPT, which is then used to generate the AI service report, as illustrated in Task 4 of Fig. 3. It is worth noting that the generated report can be tailored to different audiences, such as end users, regulators, or developers, by making minor adjustments to the prompting strategy, thereby enabling context-specific interpretations and supporting diverse analytical needs.
2.5. Experiment Configuration
In this study, to instruct ChatGPT to execute each sub-step precisely, we carefully design natural language prompts, which are presented in Fig. 3. To further guide ChatGPT in understanding the intended analysis tasks, we provide five input–output examples to enable few-shot in-context learning(brown2020language; wang2020generalizing). In our implementation of LLMAID, we utilize the default 4K-token context window provided by the ChatGPT API. However, due to ChatGPT’s rate limit, processing one app with one component per request is both time-consuming and inefficient. Conversely, submitting too many components in a single request risks output truncation or omission, which may cause misalignment between inputs and outputs. To balance task accuracy and processing efficiency, we adopt a batch size of three components per request and instruct ChatGPT to return structured outputs in JSON format. The model parameters are left at default values, except for the temperature and TopP settings, which are adjusted from 1.0 to 0.2 and 0.95, respectively. These adjustments reduce output variability while retaining a certain degree of creativity, thereby improving the consistency and reliability of the component analyses and the resulting AI service reports.
3. Evaluation
In this section, we evaluate the effectiveness of our proposed approach, LLMAID, by addressing the following research questions:
  • RQ1: To what extent can LLMAID reliably and accurately determine AI capabilities in previously identified AI Apps?
    This research question aims to assess the accuracy and reliability of LLMAID in analyzing AI capabilities within previously recognized AI Android apps. Addressing this question validates the fundamental effectiveness of LLMAID and provides the foundation for subsequent evaluations.
  • RQ2: How does LLMAID compare to the state-of-the-art rule-based approach in reliably identifying AI capabilities within in-wild Android Apps?
    We compare the performance of LLMAID with that of the SOTA rule-based method, i.e., AI Discriminator, in identifying AI capabilities across large-scale Android apps. By answering this research question, we demonstrate the effectiveness of LLMAID in accurately determining AI capabilities.
  • RQ3: How is the usefulness of the AI service summary generated by LLMAID?
    We explore the clarity and informativeness of the AI service summaries generated by our framework. These summaries are intended to provide end users, developers, and regulators with an interpretable overview of the AI functionalities embedded within a given Android application.
  • RQ4: To what extent can LLMAID uncover prevalent AI capabilities and usage patterns in real-world Android Apps?
    This final research question investigates the broader landscape of AI adoption in mobile apps by analyzing the types and distributions of AI tasks and domains identified by LLMAID. The findings offer a deeper understanding of the AI functionalities embedded in these apps and provide valuable insights to guide AI researchers and developers in designing and improving their app features.
3.1. Dataset
We perform our experiments on two separate datasets. The first comes from the existing literature, which was collected from AndroZoo (allix2016androzoo) and contains 56,682 AI Apps determined by a rule-based approach. The other is self-curated, comprising 15,517 Android apps, and reflects an open-world setting to address the absence of newer released applications, particularly those potentially integrating with emerging LLMs such as ChatGPT. Detailed descriptions of these two datasets are provided in the following paragraphs.
3.1.1. Dataset 1
Li et al. (li2022ai) proposed a SOTA rule-based approach, AI Discriminator, to identify AI-driven Android apps via keywords matching on the dissected and decompiled Android apps. They successfully identified 56,682 AI apps from a total of 7,259,232222The latest Android version is considered while multiple versions are stored. apps on AndroZoo. In our experiments, we leverage this dataset of 56,682 AI applications derived via the AI Discriminator approach to conduct cross-validation of our method, LLMAID.
3.1.2. Dataset 2
To curate our dataset, we leverage AndroZoo and extract the latest release of each Android app using metadata from its associated CSV file. For each app, we utilize the google-play-scraper npm package to verify its current availability on Google Play Store by querying with the package name (playscraper:2023), and collect its key metadata, particularly the app description and release date. In total, we identified 2,298,423 actively maintained Android apps out of a total of 23,321,823 apps available in AndroZoo at the time of our study. To focus our evaluation on apps with likely AI functionality, we further refined the dataset by examining the app descriptions retrieved from Google Play, selecting only those that strongly indicated the presence of AI-related features. Dataset 2 complements Dataset 1 by capturing newer Android apps with potential AI functionalities beyond rule-based detection, enabling a more rigorous and forward-looking assessment of LLMAID in an open-world setting.
Existing studies have often relied on manual inspection to identify AI-enabled applications; however, such an approach is inherently time-consuming and difficult to scale to the size of popular app repositories. In practice, publicly available Android applications typically include short descriptions that summarize their core functionalities, offering a lightweight entry point for automated analysis. For example, the description of ChatGO - AI Chatbot Assistant [^3] https://play.google.com/store/apps/details?id=com.chatgpt.aichat.gpt3.aichatbot mentions that ‘Discover ChatGO - AI Chatbot Assistant, the optimized [AI assistant] that transforms the way you work, create, and connect! Built on cutting-edge technologies … writing, designing, summarizing, or just exploring new ideas, ChatGO is your [go-to AI companion] … and highly personalized assistance’. This example illustrates that AI-related applications include explicit references to AI-related terms (as highlighted in red) within their descriptions. Motivated by this observation, we design a two-stage process for identifying AI applications. In the first stage, a rule-based filter is applied to detect potential AI apps based on the presence of AI-related keywords in their descriptions. In the second stage, we employ ChatGPT to perform a deeper semantic analysis of the filtered descriptions, thereby verifying whether the applications are likely to incorporate AI technologies beyond superficial references to AI-related keywords.
Specifically, the first filtering stage applies rule-based keyword extraction to identify potential AI applications based on their descriptions. The keyword set covers a broad spectrum, ranging from high-level terms (e.g., “AI” and “Deep Learning”) to more task-specific ones (e.g., “text classification” and “facial recognition”). By performing exact matches against this vocabulary, the filter effectively narrows the candidate pool, enabling subsequent analysis to concentrate on apps with higher potential relevance. However, the mere occurrence of such keywords does not necessarily imply that AI constitutes a core functionality of the app. To address this limitation, we introduce a second filtering stage powered by ChatGPT, which interprets the semantic context of app descriptions to assess whether AI techniques are likely to be employed. This stage improves the precision of the identification process by moving beyond simple keyword occurrence matching. Through this two-stage design, our approach mitigates the limitations of purely rule-based methods and more accurately detects AI-enabled Android applications. Nonetheless, due to semantic inflation and the frequent misuse of AI-related terms (e.g., “deep learning”), even when combining keyword matching with semantic analysis of descriptions, reliance on literal text alone remains inadequate and can lead to inaccurate identification. Therefore, this dataset is treated as an open-world setting and is used to evaluate the proposed approach, where incorporating implementation details enables more reliable detection of true AI apps.
Figure 4. The distribution of collected apps over the years.
In total, we successfully identified 15,517 Android applications (considering only the latest released version) from a corpus of 23,321,823 apps available on AndroZoo. Furthermore, we extracted release dates from the metadata provided on the Google Play Store and analyzed the temporal distribution of these apps, as shown in Fig. 4. The results indicate that the number of released apps has grown in parallel with the increasing prevalence of AI technologies over the past decade. Notably, Fig. 4 presents only 15,055 applications, as release date metadata was unavailable for 462 apps, accounting for 2.98% (462 out of 15,517) of the total. This minor omission does not affect the overall observation, and the upward trend of AI technology adoption in mobile applications over the past decade remains evident.
3.2. RQ1: To what extent can LLMAID reliably and accurately determine AI capabilities in previously identified AI Apps?
In this research question, we examine whether LLMAID can effectively identify AI capabilities within a set of pre-labeled AI applications. For this purpose, we employ Dataset 1 as our evaluation benchmark. Since Dataset 1 identifies AI apps but does not provide ground-truth labels for their underlying AI components, we randomly sample a subset of 382 applications and conduct human validation to assess the accuracy of LLMAID. The sample size is calculated based on the popular online Sample Size Calculator [^4] https://www.surveysystem.com/sscalc.htm with a confidence level of 95% and a margin of error of 5%.
We first present the quantity distribution of three representative types of components, including packages, APIs, and HTTPS requests in Fig. 5. Specifically, Fig. 5 shows the total number of packages, APIs, and HTTPS requests after completing Task 1 (Candidate Extraction) of LLMAID, while Fig. 5 presents the identified AI-related packages, APIs, and HTTPS requests after finishing Task 3 (AI capability analysis and detection) of LLMAID. The median values for the total considered packages, APIs, and HTTPS requests in Fig. 5 are 433.5, 296, and 53.5, respectively. It is also worth noting that the considered packages, APIs, and HTTPS requests are proactively filtered results obtained by applying a self-curated whitelist of commonly observed non-AI components during the candidate extraction process (i.e., Task 1). The whitelist of non-AI components comprises the official Java and Android kernel-related packages and API signatures. After performing AI capability analysis and determination, the median number of AI-related packages, APIs, and HTTPS requests drops to 9.5, 33, and 0, respectively. Notably, packages are filtered more aggressively than APIs, which is expected given that many AI-specific APIs are concentrated within a smaller set of AI-related packages. In contrast, HTTPS requests are sparsely detected, primarily due to the relatively recent adoption of AI-related network calls and their limited presence in the collected dataset.
(a) Considered.
(b) Determined.
Figure 5. The determination of AI candidates.
To evaluate the reliability of LLMAID in identifying AI components within the sampled dataset, we conducted a manual annotation study involving two independent authors, who independently labeled a total of 49,371 components extracted from the 382 sampled AI apps. When disagreements arise, dedicated meetings with an additional author are held to facilitate consensus. Each component was classified as either an AI component or a non-AI component. We defined a True Positive as a component correctly identified by LLMAID as AI-related, while a True Negative referred to a component correctly identified as non-AI. Following human validation, LLMAID achieved a precision of 98.05% and a recall of 93.31%, indicating strong performance in correctly detecting AI components. Furthermore, the inter-annotator agreement, measured using Cohen’s unweighted kappa coefficient (cohenkappa), was 0.82, indicating substantial consistency among the human evaluators. These results collectively demonstrate the high accuracy and robustness of LLMAID in detecting AI components.
In addition, we validated the reliability of LLMAID in identifying AI apps. We defined an app as an AI app if it contains one or more AI components determined by LLMAID. Our results showed that LLMAID correctly identified 378 out of 382 sampled applications as AI apps, corresponding to an accuracy of 99.0% (378/382). For the remaining four applications, we further manually inspected them and observed that the only potential AI indicators were filenames with the .model suffix. Without further complementary information, such as associated APIs, packages, or functional descriptions, LLMAID was unable to classify these apps as AI apps, hence excluding them from the set of AI apps.
Table 1. Comparison of the reliability of LLMAID between the implementations of analysis-then-detection and detection-then-analysis.
Analysis-then-Detection Detection-then-Analysis
AI Candidate
AI Apps
Time (sec)1
AI Candidate
AI Apps
Time (sec)
2859 378 638.76 2192 371 226.94
  • 1
    Average running time for an Android APK in seconds.
Although the design of Task 3.1 (AI Capability Analysis) and Task 3.2 (AI Component Determination) is inspired by Chain-of-Thought prompt (wei2022chain), this sequence may not be optimal for LLMAID. A more efficient alternative is to first determine whether a component is AI-related or not and then analyze and interpret only those classified as AI components, effectively reversing the order of Tasks 3.1 and 3.2. Under the same evaluation setup on the 382 sampled apps, we find that this detection-then-analysis approach reduces the average processing time per app by 65%, as summarized in Table 1. However, the number of AI components identified using this approach is notably lower than that obtained with the original sequential implementation (Analysis-then-Detection). Based on the overall results, we suggest that detection-then-analysis can only be used when developers are interested in finding AI apps. If developers seek to understand all possible AI capabilities of an app, we foresee that the original LLMAID with analysis-then-detection is more effective in capturing AI candidates.
RQ1 Findings Human validation suggests that LLMAID achieves high precision (98.05%) and recall (93.31%) in detecting AI components and AI apps. Although reversing the sequence of Task 3.1 and Task 3.2 could improve the efficiency of LLMAID, it compromises reliability compared to the original sequential implementation.
Table 2. Number of AI and Non-AI apps classified by AI Discriminator and LLMAID.
AI Discriminator LLMAID
AI Non-AI AI Non-AI
555 3264 1366 2453
Table 3. Detected AI Apps by AI Discriminator and LLMAID
AI Discriminator
AI Discriminator
LLMAID
LLMAID
9 546 820
3.3. RQ2: How does LLMAID compare to a state-of-the-art rule-based approach in reliably identifying AI capabilities within in-wild Android Apps?
In this research question, we evaluate how reliable LLMAID is in determining AI capabilities and compare it to the SOTA approach, AI Discriminator (li2022ai; aiapp:2023). To efficiently perform the evaluation, we select a representative subset of Dataset 2, focusing on the apps published in the recent five years, and randomly sampling 30% apps in each year to preserve the same quantity distribution. Finally, we ended up with 3,819 $(1980\times 30\%+2336\times 30\%+2326\times 30\%+3053\times 30\%+3044\times 30\%)$ apps for the experiment.
Following the same experimental setup as in RQ1, we compute the precision and recall for AI component identification. LLMAID achieves a precision of 97.32% and a recall of 91.01%, demonstrating strong performance in identifying AI components. The inter-annotator reliability, measured using Cohen’s unweighted kappa coefficient, is 0.80, which reflects a high level of agreement with the results generated by LLMAID.
Table 3.2 presents the experimental results for AI App identification, comparing the performance of AI Discriminator with that of LLMAID. Among the total of 3,819 experimental Android Apps, 555 of them are classified by AI Discriminator as AI Apps, while 1,366 are recognized as AI Apps by LLMAID, which is 242% more than AI Discriminator. Table 3.2 shows the number of AI Apps identified by AI Discriminator and LLMAID. It clearly demonstrates that 546 Android Apps, accounting for 90.55% $(546/603)$ of those identified by AI Discriminator and 39.97% $(546/1366)$ of those confirmed by LLMAID, are commonly recognized as AI Apps by both approaches. To manually inspect the nine AI apps identified solely by AI Discriminator, multiple authors downloaded the APK files from AndroZoo and dissected them with Jadx (jadx2023). Two apps are recognized as AI apps because only one special tflite file is being detected, while the remaining seven apps are classified as AI apps because of keywords, such as “DecisionTreeClassifier” and “KNeighborsClassifier” in their descriptions of json files. However, the description contents in the JSON files alone, without supporting code-level detection, may not imply the actual implementation of AI capabilities and therefore should not be regarded as convincing AI evidence.
To verify the 820 apps only classified by LLMAID, we randomly sampled 86 apps (with a confidence level of 95% and a margin of error of 10%) and cross-validated them using the decompiler tool, Jadx (jadx2023), with multiple authors involved. In the course of this cross-validation, we observed that OpenAI APIs were sometimes directly invoked within Java code, whereas in other cases developers implemented these API calls through native libraries (i.e.,.so files). Such implementations of cloud-based AI capabilities could not be detected by AI Discriminator due to the limitations of its rule set. In the end, we confirmed that 83 of the total of 86 sampled apps are classified by LLMAID as truly AI apps, showcasing a high level of identification precision.
RQ2 Findings High precision (97.32%) and recall (91.01%), and the high inter-annotator reliability about the identification of AI components reconfirm the effectiveness of LLMAID for in-wild apps. Furthermore, the experimental result of detecting more than twice as many AI apps as the rule-based approach instantiates the outperformance of LLMAID against the SOTA one.
Figure 6. Description of Copy.ai (the left) and its AI service summarization (the right) generated by our framework.
3.4. RQ3: How is the usefulness of the AI service summary generated by LLMAID?
In this research question, we seek to evaluate the usefulness of the AI summary, which is generated in Task 4 of LLMAID based on the app’s implementation details. Specifically, we compare the generated AI summaries with the developer-provided descriptions on the app store, using this comparison as a diagnostic tool to reveal the shortcomings of the developer-provided information and to enable a more accurate understanding and evaluation of AI functionalities.
We conducted a user study by recruiting six developers, each with at least three years of experience in Android app analysis. Specifically, we randomly selected 20 Android apps identified as AI apps in RQ2, and then generated their AI summaries. For each app, we paired each LLMAID generated AI summary with its corresponding app description retrieved from the Google Play store via the google-play-scraper (playscraper:2023), and displayed them side-by-side to participants, as shown in Fig. 6. We divided the six participants as two groups. Each group evaluated ten Android apps and each app was evaluated by three participants individually. Before conducting the user study, participants were instructed to download the apps and get familiar with their major features. We then asked participants to read and compare the app description and the AI summary for each app, evaluating both from five aspects: Accuracy (i.e., Do you think that the summary is accurate concerning the AI functions the app provides?), Completeness (i.e., Do you think that the summary is complete about the provided AI tasks?), Consistency (i.e., Do you think that the summary is concrete enough concerning the AI functions?), and Recommendations for app users (i.e., Would you like to recommend the summary to other app users?) and developers (i.e., Would you like to recommend the summary to other app developers to inspire new app development?). Each dimension was assessed using a 10-point Likert scale, with scores ranging from 1 (indicating poor quality) to 10 (indicating high quality).
Table 4. Average and median scale values voted for sampled Android apps by six participants. “Recom.” stands for Recommendation, “devs” stands for developers, “Med.” stands for median.
Summary Type Accuracy Completeness Consistency Recom. app users Recom. app devs.
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
[…]
Table 4 shows the user study results measured by the mean and median across five dimensions. It shows that the AI summary generated by LLMAID consistently outperformed the original app descriptions, underlining by higher mean scores (+2) and median scores (+1). In terms of Accuracy and Completeness, the AI summary attained mean scores of 8.67 and 8.63, respectively. These results underscore the effectiveness and robustness of the proposed AI detection framework in capturing and conveying relevant AI functionalities for mobile apps. Moreover, the standard error of the mean of the summary generated by LLMAID is approximately one-third of the original description of the app, indicating a greater consensus among participants on the quality of the summaries generated by LLMAID. However, the voted scores for both the summary generated by LLMAID and the original app description are quite close with respect to the metric of Recommendation for app users. Specifically, the median value is the same, and the difference between the average scores is only 0.63, showcasing that different participants may have different preferences when recommending summaries to app users. In general, the generated summaries by LLMAID exhibiting higher accuracy, completeness, and consistency are well-received by developers and researchers; however, they may be less appreciated by downstream app users. To gain more valuable insights from app users’ perspective, we intend to expand the user study to common app users as future work.
RQ3 Findings The user study shows that the AI service summary generated by LLMAID is constantly preferred by practitioners in terms of accuracy, completeness, consistency, and recommendation.
Table 5. Different types of identified AI components.
Data Type
Package API
Model
HTTP Requests
Others Total
#. items 10,285 644 120 32 11,081
3.5. RQ4: To what extent can LLMAID uncover prevalent AI capabilities and usage patterns in real-world Android Apps?
Regarding this research question, we seek to classify the specific domains and tasks of the identified AI components in the determined AI apps in RQ2. Table 5 shows the number of AI components in different categories. The majority of the AI components (92.82%, 10285/11081) lie within packages and their corresponding APIs. The second largest category of AI components is model, which refers to on-device AI model files, such as .tflite and .caffemodel. Apart from package names, APIs, and on-device AI models, we observe that some apps rely on HTTP requests (e.g., OpenAI APIs) to access cloud-based AI services, which has rarely been discussed in previous studies (xu2019first; li2022ai). The remaining components exhibiting limited shared patterns are collectively classified under the category of Others.
Figure 7. The distribution of tasks (inner ring) and their corresponding AI domains (outer ring) for the AI components.
With these identified AI components, we first reviewed the analysis results to design a classification scheme that captures a wide range of AI domains and tasks. Due to the large number of AI components, we employ ChatGPT to automatically categorize each AI component into a corresponding domain and task. Fig. 7 presents the distributions of classified domains and tasks of the determined AI components. The area of each colored sector represents its corresponding proportion. At the top level, AI components are classified into six common AI domains, including Computer Vision (54.80%), Data Analysis (26.85%)  [^5] Data analysis and processing denote statistical operations within AI frameworks., Natural Language Processing (13.38%), Others (2.75%), Audio and Speech Processing (1.90%), and Augmented Reality (0.32%). The category Others contains AI components that are difficult to classify clearly. The results show that the dominant AI capabilities are still oriented toward Computer Vision and most AI components target the tasks of Object Detection (45.97%), Text Recognition (17.95%), and Face Recognition (16.77%). The second largest AI domain is Data Analysis, which is occupied mostly by the task of Data Processing (86.69%), such as tensor operations and input/output data parsing.
Table 6. Number of Android apps in different domains.
Domains
Data Analysis
Computer Vision
Natural Language
Processing (NLP)
Audio & Speech
Processing (ASP)
Augmented Reality
Others
#. Apps 643 634 43 30 2 14
Based on classified AI components, we further categorize the domains and tasks at the app level. For each application, we begin by tallying the occurrence of domains and tasks associated with its AI components, and then categorize the app under the domain and task with the highest frequency. Table 6 shows the number of apps in different domains. The majority of AI apps fall into the Data Analysis domain, whereas most AI components are categorized under Computer Vision. This discrepancy is primarily due to the reuse of common components across various apps, with a significant concentration in the Computer Vision category. In addition, the number of apps in both the Computer Vision and Data Analysis domains is nearly equal, making them the two most prominent categories in both AI component classification and app-level categorization.
In summary, the majority of AI components predominantly utilize machine learning and data processing algorithms to offer vision-related features, aiding in everyday tasks like face recognition and handwritten character recognition. In contrast, virtual reality functions are minimally incorporated in these apps. This could be due to the fact that virtual reality functions may lack compelling application scenarios and are not currently considered essential by developers and users.
RQ4 Findings The most widely used AI services among Android apps and their underlying AI components are Computer Vision (54.80%) and Data Processing (26.85%), respectively. This may be attributed to the widespread user demands for vision-related functionalities, which are commonly supported across various types of mobile apps. In contrast, the augmented reality oriented ones are the least provided due to the limited number of compelling real-world application scenarios, despite the concept’s novelty and growing popularity.
4. Discussion
We now discuss the inherent limitations (Section  4.1) and possible implications for both researchers and practitioners based on the outcomes of our study (Section 4.2).
4.1. Limitations
4.1.1. LLM Hallucination
Hallucination, referring to incorrect or fabricated responses to user inquiries, and inconsistency in responding to identical queries are inherent limitations of LLMs. In this study, we inquire about ChatGPT with our extracted implementation details (i.e., components) of Android apps to determine whether they are AI-related or not. However, the responses given by ChatGPT may not be as accurate as they really are due to hallucinations and output inconsistency. To mitigate such threats, we employ carefully crafted prompts and leverage large-scale experimental datasets to reduce the likelihood of hallucinated outputs and enhance response consistency. Our manual inspection of the sampled experimental results indicates that LLMAID delivers reliable and accurate analyses, demonstrating its robustness despite the inherent limitations of LLMs.
Figure 8. The percentage of obfuscated Java file names and their directory names.
4.1.2. Obfuscations in Android apps
Obfuscation techniques, such as string encryption (e.g., the obfuscated API: <<d.f.e.x.b: void <<init>> (int)>>) and Java reflection, are widely adopted in Android apps to safeguard intellectual property against reverse engineering, as extensively evidenced by existing research (dong2018understanding; glanz2020hidden). We have attempted to replicate some deobfuscation artifacts, such as StringHound (glanz2020hidden) and simplify (guillot2010automatic), to recover the obfuscated items in the Android apps in our study. However, these tools either failed to execute successfully in our environment or produced unsatisfactory results. To evaluate how significantly obfuscations affect the final results, we measured the ratio of obfuscation in Android apps in Dataset1. Fig. 8 shows the percentage of obfuscated Java file names and the directory names they reside in. The median value of the obfuscated Java file names and their corresponding folder names account for 11.94% and 20.23%, respectively. While obfuscation may inevitably introduce some limitations, the substantial portion of unobfuscated implementation details still enables us to extract meaningful insights and supports the reliability of our overall analysis and summarization.
4.2. Implications
4.2.1. On-Device and cloud AI integration in Android apps.
The growing integration of on-device models with cloud-based LLM APIs in Android applications introduces new opportunities and challenges that warrant systematic investigation. While such hybrid architectures promise enhanced proactivity, context-awareness, and personalization, they also raise unresolved issues of reliability, privacy, and ethical responsibility. Future research should extend LLMAID to analyze how Android apps orchestrate local and cloud intelligence, assess resilience under failures or connectivity disruptions, and examine the privacy–utility trade-offs inherent in offloading user data to remote services. Equally important is understanding the evolution of developer practices in adopting hybrid AI features, from SDK/API usage to prompt design and deployment strategies, as well as addressing transparency and fairness concerns in user-facing AI behaviors. Advancing these directions will be essential for establishing trustworthy, adaptive, and ethically responsible AI integration in the Android ecosystem.
4.2.2. Security mitigation for software artifacts.
Our study reveals that by asking LLM about the implementation basics, it could respond with accurate, at least to some extent, descriptions of whether they are AI-related or not. One possible promising future research direction would be to use LLM as a neural knowledge to inquire if the API involves some existing vulnerabilities, for example, the vulnerability record in public Common Vulnerabilities and Exposures (CVE). Both researchers and developers could benefit from LLM-based vulnerability inquiries in that they could efficiently and effectively respond to potential vulnerability fixes throughout the software engineering process. Consequently, more reliable and secure artifacts could be achieved.
4.2.3. Software artifacts analysis for researchers.
Our research on the summarization of software functionality in RQ3 validates the potential of LLM in code interpretation and summarization, specifically within the context of Java as a single programming language. The widespread use of cross-platform frameworks necessitates code interpretation and summarization across multiple programming languages, as these frameworks often integrate components written in different languages to ensure compatibility across platforms. However, it is not straightforward to merge the interpretation in different programming languages due to the interactions between them. Detailed analysis of interactions between different languages and careful design of prompts are necessary to make full use of the powerful potential offered by LLMs, which also paves a promising research direction for our research communities.
4.2.4. Advanced AI identification approaches & questionnaire with app developers.
In the AI app identification process, we found that some self-claimed AI Android apps, despite describing advanced AI features in their Play Store listings, contained no detectable AI components in the released APKs. A possible reason is the use of Android App Bundles (appbundle:2023), which deliver only essential components at the installation time and provide the necessary AI functions in the on-demand parts, preventing full inspection of the implementation details. To enable comprehensive analysis, it is imperative to adopt a new approach capable of harvesting all parts of an Android app, not only the basic ones but also those loaded on demand. Moreover, not all AI function implementations in Android apps are trivial. There may be some non-trivial implementations that are transparent to us. Therefore, a questionnaire with these popular Android app developers would provide us with a few great insights for further app analysis.
5. Related Work
5.1. AI Mobile Apps Study
Xu et al. (xu2019first) first proposed an APK analysis tool, DL Sniffer to empirically investigate the use of deep learning techniques in Android apps. The DL Sniffer dissects the released APK and identifies the utilization of 16 popular deep learning frameworks via two separate modules. The first one recognizes specific strings (e.g., “TF_AllocateTensor”) in the section rodata of native shared libraries (i.e., file ends with “.so”) via some powerful Linux command (readelf:2023; objdump:2023). The other one identifies AI special APIs in DEX binary files by de-compiling the binary file into smali code via (dex2jar:2023). Experiments on the large-scale dataset containing 16,500 Android apps reveal that top apps are AI techniques early adopters, especially photo beauty apps, and the on-device AI is actually lighter than practitioners expect. Sun et al. (sun2021mind) proposed ModelXRay to determine whether on-device AI frameworks are adopted and to check if the adopted framework is protected or encrypted in a large dataset containing 46,753 Android apps. Their experimental results show that 41% Android apps do not protect the adopted DL models, and even some of the protected models still have a weak status, which would result in the leakage of private information and economic losses. Li et al. (li2022ai) conducted an exploratory study and released a tool, AI Discriminator, to analyze the adoption of AI technologies in Android apps. They first refined the rules used in the identification of AI app and successfully recognized 56,682 AI apps from a total of 7,259,232 mobile apps provided on AndroZoo (allix2016androzoo). With the historical versions of apps considered, they concluded that the AI models utilized in Android apps are not updated regularly and are not protected carefully. What’s more, AI apps always request access to various types of sensitive user data. In addition, they analyzed reviews for the determined AI apps and revealed that users always express positive attitudes towards the adoption of advanced AI technologies.
5.2. LLMs in Software Engineering
Feng et al. (feng2023prompting) tried to reproduce Android bugs with the help of LLMs. They proposed a lightweight bug replay tool, AdbGPT, to automatically rerun Android APKs and trigger bugs according to collected bug reports. The AdbGPT first utilizes prompt engineering , working as real app developers, to extract the necessary steps for bug replay from the given bug reports. And then prompts LLMs to run Android apps under the guidance extracted previously. Their extensive experiments reveal that AdbGPT could obtain a higher accuracy in necessary step extraction compared with the state-of-the-art traditional approaches and finally successfully reproduce 81.3% bugs, which is much better than baselines. Their approach demonstrates the effectiveness of LLMs in assisting traditional software research problems. Tian et al. (tian2023chatgpt) focused on the research tasks of code generation, program repair, and code summarization and conducted an exploratory study to investigate performance in the aforementioned practices. They curated datasets from LeetCode 2016-2020 and LeetCode 2022, and constructed prompts to generate code with problem descriptions, repair buggy code implementations, and summarize code intentions. The experimental results not only reveal that ChatGPT is effective in common software practices but also unveil that limitations exist in LLMs, such as long and detailed prompt descriptions that would hinder the effectiveness of ChatGPT. Ma et al. (ma2023scope) conducted a study to touch on the limitations of the interpretability of ChatGPT in three types of software engineering jobs, including understanding syntax, static behaviors, and dynamic behaviors. Performing ChatGPT on different programming languages, the authors concluded that ChatGPT does have the ability to construct AST but it still produces incorrect or fake facts with respect to dynamic semantic understanding. It is necessary for researchers to propose effective methods to verify the correctness of the generated outputs from ChatGPT. Li et al. (li2023hitchhiker) carried out an LLM-assisted static analysis with respect to use-before-initialization (UBI) bugs for C/C++ programs. They proposed an automatic agent that interacts with both a static analysis tool and an LLM. With the state-of-the-art static analysis tool and meticulous design of prompts, the authors could successfully overcome several challenges, such as modeling bugs specific to the problem, which are hard to manage within separate research directions. They conducted experiments on real potential UBI bugs, achieved higher precision and recall, and even identified 13 previously untouched ones, revealing the effectiveness of LLM in assisting software engineering problems.
6. Conclusion
In this paper, we have proposed an automated approach LLMAID, which employs program analysis techniques to extract detailed implementation components from Android apps and utilizes a large language model (ChatGPT in this study) to interpret and summarize the underlying AI services. Specifically, our approach addresses the intuitive questions, including “Does the application incorporate AI capabilities” and “What specific types of AI functionalities are embedded”. The experimental results on both AI apps previously identified by a rule-based approach and in-wild apps demonstrate that LLMAID could achieve a precision of 98.05% and a recall of 93.31% on the former AI apps and a precision of 97.32% and a recall of 91.01% on the latter in-wild apps regarding AI components identification. Moreover, LLMAID not only identifies more than twice as many AI apps in the wild compared to the SOTA rule-based approach, but also provides meaningful functional descriptions. To further evaluate its practical utility, we have also conducted a user study on the usefulness of LLMAID and the study results confirm that the app summary generated by LLMAID is much preferred by participants with respect to the metrics of accuracy, completeness, consistency, and recommendation. Additionally, our analyses reveal that the most prevalent AI services are related to machine learning (26.85%) and computer vision (54.80%), whereas virtual reality based functions are actually the least presented, despite their growing popularity and increasing interest among both AI researchers and early adopters. Furthermore, with these findings, we outline several promising research directions to inform and inspire future work in this area.
Data Availability: The source code and datasets are all made publicly available in our artifact package via the following link: https://zenodo.org/records/15522834

### v2

@@
-LLMAID: Identifying AI Capabilities in Android Apps with LLMs
-Abstract.
+
+Abstract
 Recent advancements in artificial intelligence (AI) and its widespread integration into mobile software applications have received significant attention, highlighting the growing prominence of AI capabilities in modern software systems.
@@
 We apply LLMAID to a dataset of 4,201 Android applications and demonstrate that it identifies 242% more real-world AI apps than state-of-the-art rule-based approaches.
-Our experiments show that LLM4AID achieves high precision and recall, both exceeding 90%, in detecting AI-related components.
+Our experiments show that LLMAID achieves high precision and recall, both exceeding 90%, in detecting AI-related components.
 Additionally, a user study indicates that developers find the AI service summaries generated by LLMAID to be more informative and preferable to the original app descriptions.
@@
 The results reveal a strong concentration of AI functionality in computer vision (54.80%), with object detection emerging as the most common task (25.19%).
-1. Introduction
+1 Introduction
 Advancements in artificial intelligence (AI) and their successful integration into mobile applications (apps) have enhanced the capabilities of Android apps in daily activities, such as optical character recognition, autonomous driving, object recognition and shopping recommendations.
-To support developers’ integration of AI capabilities, several vendors have introduced mobile AI frameworks, such as TensorFlow Lite (abadi2016tensorflow), Caffe2 (caffe2:2023), Core ML (coreml:2023), and ncnn (ncnn:2023).
-These frameworks enable efficient on-device model building and inference, offering better privacy protection and reducing bandwidth usage compared to cloud-based models (guo2018cloud), while also improving responsiveness and ensuring functionality in resource-constrained environments typical of mobile devices.
-In recent years, there has been a surge in the development of large AI models, such as stable diffusion (stable:2023), ChatGPT (chatgpt:2023), LLaMA (touvron2023llama), and DeepSeek (liu2024deepseek).
-Unlike traditional lightweight frameworks, these models provide more advanced capabilities in reasoning, multimodal content understanding, and generative tasks.
-With their advanced abilities to understand and generate content, emerging app developers increasingly resort to these models via the cloud to enhance their businesses (xu2019first; hu2023first), enabling the creation of richer user experiences such as personalized recommendations, interactive assistants, and multimodal content generation.
+To support developers’ integration of AI capabilities, several vendors have introduced mobile AI frameworks, such as TensorFlow Lite abadi2016tensorflow, Caffe2 caffe2:2023, Core ML coreml:2023, and ncnn ncnn:2023.
+These frameworks enable efficient on-device model building and inference, offering better privacy protection and reducing bandwidth usage compared to cloud-based models guo2018cloud.
+In recent years, there has been a surge in the development of large AI models, such as stable diffusion stable:2023, ChatGPT chatgpt:2023, LLaMA touvron2023llama, and DeepSeek liu2024deepseek.
+With their advanced abilities to understand and generate content, emerging app developers increasingly resort to these models via the cloud to enhance their businesses xu2019first; hu2023first.
 With the widespread adoption of AI and the growing awareness of its inherent limitations, issues of AI discoverability and transparency have attracted attention from both application consumers and various regulatory bodies.
@@
 In this paper, we define this process as AI Discovery, which encompasses determining whether AI capabilities are integrated or not and further elucidating the functional characteristics of the incorporated AI services.
-Several studies have been conducted to investigate AI techniques in smartphone apps (xu2019first; zhao2022survey; li2022ai; siu2023towards; deng2019deep).
-Specifically, Deng et al.
-(deng2019deep) and Zhao et al.
-(zhao2022survey) both provided comprehensive surveys on deep learning for mobile devices, highlighting hardware architectures, software frameworks, and optimization mechanisms such as pruning, quantization, and hardware accelerators that enable efficient, privacy-preserving, and cost-effective on-device AI across diverse application domains.
-Siu et al.
-(siu2023towards) empirically investigated user–AI interaction in 176 mobile apps, identifying 255 AI features and 759 implementations, ultimately categorizing interaction patterns and demonstrating the importance of usable interaction design for on-device AI features.
-Li et al.
-(li2022ai) conducted the largest empirical study to date, analyzing 56,682 AI apps with their newly proposed rule-based approach, analyzing application trends, framework adoption, model protection, and user-related concerns, and releasing an open dataset for future research.
-Xu et al.
-(xu2019first) examined 16,500 popular Android apps to reveal early adopters of deep learning, their usage scenarios, and model characteristics, offering insights that emphasize both the potential of mobile deep learning and the need for optimizations, model protection, and validation.
-However, these studies exhibit several notable constraints: 1) lacking attention to other potential AI components, [^1] An AI component may refer to either high-level modules (e.g., task planning engines) or low-level implementation artifacts (e.g., APIs, libraries, HTTP requests), whereas an AI feature denotes high-level functionalities (e.g., image recognition, sentiment analysis) enabled by these components.
+Several studies have been conducted to investigate AI techniques in smartphone apps xu2019first; zhao2022survey; li2022ai; siu2023towards; deng2019deep.
+However, these studies exhibit several notable constraints: 1) lacking attention to other potential AI components, [^1] Component in the paper may refer to either high-level modules or low-level elements such as APIs and libraries, depending on context.
 such as integrated development packages, internal assets, and HTTP requests and API invocations in detailed implementations and integrated libraries.
@@
 To address these aforementioned challenges, we argue that automating the human-like semantic-based AI Discovery is essential, enabling the characterization of real-world AI components by interpreting their meaningful names.
-Inspired by recent findings (hagendorff2023human; ning2024can; li2024llms; pan2024large) that large language models (LLMs) are capable of exhibiting human-like reasoning and decision-making, we explore how these models can enhance the AI Discovery process by interpreting semantically meaningful AI component names.
-Moreover, recent studies have shown that these LLMs can automate various software tasks, including automatic program repair (zhang2022repairing; jiang2023impact; xia2023conversational; han2024chase), code generation (nijkamp2022conversational; chakraborty2022natgen; fried2022incoder; liao20243), and improving software analysis efficiency and effectiveness (ma2023scope; li2023hitchhiker).
+Inspired by recent findings hagendorff2023human; ning2024can; li2024llms; pan2024large that large language models (LLMs) are capable of exhibiting human-like reasoning and decision-making, we explore how these models can enhance the AI Discovery process by interpreting semantically meaningful AI component names.
+Moreover, recent studies have shown that these LLMs can automate various software tasks, including automatic program repair zhang2022repairing; jiang2023impact; xia2023conversational; han2024chase, code generation nijkamp2022conversational; chakraborty2022natgen; fried2022incoder; liao20243, and improving software analysis efficiency and effectiveness ma2023scope; li2023hitchhiker.
 These findings provide evidence that LLMs may have substantial knowledge in software practice, motivating us to explore their full potential in AI Discovery within mobile applications.
@@
 This structured approach allows us to perform AI discovery in an automated, step-by-step manner, significantly reducing the need for manual intervention.
-We implemented LLMAID using the widely adopted GPT-3.5-turbo model, which is much more cost-effective than the GPT-4 model (gptmodel:2023).
-To evaluate the effectiveness of LLMAID, we conducted validation on a sample of 382 AI applications selected from a total of 56,682 already identified AI apps (li2022ai).
+We implemented LLMAID using the widely adopted GPT-3.5-turbo model, which is much more cost-effective than the advanced GPT-4 model gptmodel:2023.
+To evaluate the effectiveness of LLMAID, we conducted validation on a sample of 382 AI applications selected from a total of 56,682 already identified AI apps li2022ai.
 Manual evaluation of the results confirms that LLMAID achieves high precision and recall (exceeding 90%) in detecting AI components.
 Furthermore, we applied LLMAID to a dataset of 3,819 real-world Android apps published on Google Play between 2019 and 2023.
-Compared to the state-of-the-art (SOTA) detection approach, AI Discriminator (li2022ai), LLMAID identifies more than twice as many AI Android apps, showcasing its superior detection capability.
+Compared to the state-of-the-art (SOTA) detection approach, AI Discriminator li2022ai, LLMAID identifies more than twice as many AI Android apps, showcasing its superior detection capability.
 Our analysis reveals that the majority of detected AI components pertain to the domains of computer vision (54.80%) and data analysis (26.85%), with object detection (25.19%) and data processing (23.17%) being the most prevalent tasks.
@@
   • We conduct the first analysis to uncover prevalent AI domains and tasks within Android apps.
-Figure 1. The workflow of LLMAID.
-2. Approach
+Figure 1: The workflow of LLMAID.
+2 Approach
 Fig.
 1 illustrates the general framework and workflow of LLMAID (LLM for AI Discovery).
-As mentioned above, the LLMAID framework consists of four stages: candidate extraction, knowledge base interaction, AI capability analysis and detection, and AI service summarization.
-Together, these stages are designed to address the complexity of uncovering AI functionalities in Android apps.
-In the first stage, Candidate extraction, static analysis techniques are applied to the Android APK to extract fine-grained components that may potentially involve AI functionality.
-The second stage, Knowledge base interaction, refines these candidates by referencing dynamically constructed AI and non-AI knowledge bases, which act as caches of previously analyzed components.
-This step reduces redundant analysis and improves precision by filtering out irrelevant components early.
-The third stage, AI capability analysis and detection, prompts a large language model to interpret the functionality of each remaining candidate and determine whether it constitutes an AI component, yielding a curated set of AI-relevant elements together with functional descriptions.
-Finally, the AI service summarization stage produces a concise AI service report for the given APK by leveraging a carefully designed LLM prompt, which incorporates the determined AI components along with their interpreted AI functional descriptions.
-Importantly, the overall workflow is model-agnostic and can be applied with any advanced large language model, ensuring adaptability as LLM technologies continue to evolve.
-We elaborate on each stage in the following subsections.
-Figure 2. The structure of a typical released Android APK.
-2.1. Task 1: Candidate Extraction
-To determine AI capabilities within an Android application, we first extract all relevant information—referred to as candidates-from the APK file.
+As mentioned above, the LLMAID framework consists of four stages.
+These stages are designed to address the complexity of uncovering AI functionality in Android apps.
+Candidate extraction employs static analysis techniques to extract potential AI components from an Android APK.
+Knowledge base interaction aims to filter these extracted components using a constructed AI and non-AI knowledge base, thereby improving the accuracy of downstream AI capability analysis.
+The AI capability analysis and detection stage further identifies and characterizes AI components by prompting an LLM to analyze.
+Finally, AI service summarization automatically generates a report of these AI components for the given Android APK.
+Importantly, our approach is model-agnostic and can be applied to any advanced large language model.
+We describe each stage in the following subsections.
+Figure 2: The structure of a typical released Android APK.
+2.1 Task 1: Candidate Extraction
+To determine AI capabilities within an Android application, we first extract all relevant information—referred to as candidates—from the APK file.
 Fig.
-2 shows the structure of a typical released Android APK file, typically including classes.dex, which is the Dalvik executable file (dalvik:2022), assets, META-INFO, lib, AndroidManifest.xml, res, and resources.arsc.
-Notably, an APK may comprise multiple .dex files because each file is constrained by the 65,536-method limit.
-When this threshold is exceeded, the compiler automatically partitions the code into additional files, conventionally named classes1.dex, classes2.dex, and so forth.
-We utilize AAPT (aapt:2024) to extract the structure from the curated Android APK file.
+2 shows the structure of a typical released Android APK file, typically including classes.dex, which is the Dalvik executable file dalvik:2022, assets, META-INFO, lib, AndroidManifest.xml, res, and resources.arsc.
+Notably, an APK may contain multiple .dex files due to the method limit of 65,536 per file.
+We utilize AAPT aapt:2024 to extract the structure from the curated Android APK file.
 Our focus is on examining the implementation details embedded in the app, which requires decompiling .dex files to retrieve the underlying Java logic and analyzing .so files in the lib directory for native code.
 Specifically, we extend a static analysis approach, namely App Scanner, proposed by Liu et al.
-(liu2021identifying) to extract APIs from an APK file.
-The scanner leverages the Soot framework (lam2011soot), a widely used Java-based static analysis framework that has been extended to support Android applications.
-In particular, Soot translates Dalvik bytecode (dalvik:2022), the default executable format for Android applications generated from Java (or Kotlin) source code through an intermediate compilation step and optimized for devices with limited memory and power resources, into Jimple (bartel2012dexpler).
-Jimple is a three-address intermediate representation that is both human-readable and well suited for static program analysis.
-This transformation bridges the gap between low-level Dalvik instructions and a higher-level abstraction and enables a more effective analysis of Android applications.
-The App Scanner method in this study is enhanced to additionally capture fine-grained components from multiple sources within the given APK, such as .dex files, native library .so files, and the asset files.
-This also allows us to extract more indicators of AI functionality, such as HTTP requests https://api.openai.com/v1/ and descriptive package name com.google.firebase.ml.vision.objects.
-Moreover, to analyze embedded dynamic native libraries, we employ tools readelf (readelf:2023) and objdump (objdump:2023) to parse HTTPS requests, as many apps implement AI capabilities by invoking these requests via the Android NDK (ndk:2023).
-2.2. Task 2: Knowledge Base Interaction
-Given our focus on implementation-level details of AI capabilities, the extraction process produces a substantial number of fine-grained components that include both AI-relevant and non-AI functionalities.
-The presence of non-AI components introduces noise into the analysis, as they do not contribute to the understanding of AI capabilities and can impede the accurate identification of AI-related functionality.
-To reduce such noise, we first apply a rule-based filtering approach designed to exclude non-AI components, such as official Java and Android kernel-related; however, because this method operates at a coarse granularity, some non-AI components may still be incorrectly retained as AI-related.
-To address the limitations of coarse-grained rule-based filtering and further improve analysis efficiency, we introduce a knowledge base (KB) interaction step into the workflow.
-This step enables the reuse of analysis outcomes for components that have already been processed, allowing previously verified AI or non-AI components to be retrieved directly and thereby reducing redundant computation.
-To support this process, we construct two complementary KBs that function as caches rather than shortcuts for classification: an AI KB, which stores verified AI components together with their capability analyses, and a non-AI KB, which records confirmed non-AI components to facilitate pre-analysis filtering.
-After component extraction and initial filtering, the system checks each candidate against both KBs; if a match is found, the corresponding analysis and detection result can be directly retrieved, thereby bypassing further inference.
+liu2021identifying to extract APIs from an APK file.
+The scanner aims to transform Dalvik bytecode dalvik:2022 into Jimple bartel2012dexpler, a human-readable intermediate representation.
+The App Scanner method is enhanced to additionally capture fine-grained components from multiple sources within the given APK, such as .dex files, native library .so files, and the asset files.
+This also allows us to extract more indicators of AI functionality, such as HTTP requests https://api.openai.com/v1/ and descriptive package name com.google.firebase.ml.vision.o-
+bjects
+.
+Moreover, to analyze embedded dynamic native libraries, we employ tools readelf readelf:2023 and objdump objdump:2023 to parse HTTPS requests, as many apps implement AI capabilities by invoking these requests via the Android NDK ndk:2023.
+Figure 3: Prompt design and examples for LLMAID
+2.2 Task 2: Knowledge Base Interaction
+The extraction process yields a substantial number of fine-grained components capturing both AI and non-AI implementation details.
+To filter out the irrelevant non-AI components, we apply a rule-based approach following the extraction process.
+This approach aims to remove components that are irrelevant to AI functionalities; however, it operates at a coarse granularity, which may result in the retention of some non-AI components falsely classified as AI-related.
+To further refine the analysis and enhance efficiency, we incorporate a knowledge base (KB) interaction step into the workflow.
+This step enables the reuse of prior analysis results to accelerate processing and reduce redundant computation.
+Specifically, we maintain two separate KBs: an AI KB that stores previously verified AI components and their associated capability analyses, and a non-AI KB that contains confirmed non-AI components used primarily for pre-analysis filtering.
+Once the components are extracted and initially filtered, they are checked against both KBs.
+If a match is found, the corresponding analysis and detection result are retrieved directly, bypassing further computation.
 Otherwise, the components are forwarded to the next stage, Task 3: AI capability analysis and detection, whose results are then systematically incorporated into the corresponding KB for future reference and reuse.
-To ensure reliability, both KBs are continuously refined and manually audited, ensuring that cached results maintain high accuracy while significantly reducing unnecessary computation.
-2.3. Task 3: AI Capability Analysis and Detection
-This stage, the core of LLMAID, focuses on identifying which of the extracted components are AI-related and characterizing them in terms of their AI functional capabilities.
+2.3 Task 3: AI Capability Analysis and Detection
+This stage, the core of LLMAID, involves identifying and characterizing the extracted AI components.
 To ensure precision and interpretability, this task is divided into two sub-steps: AI capability analysis and AI component determination.
-The first sub-step, AI capability analysis, interprets the functional behavior of each candidate component to identify the type of functionality it provides, such as image recognition, language processing, or network communication.
-The second sub-step, AI component determination, then decides whether the interpreted component should be classified as AI-related, filtering out non-AI elements.
-After the knowledge base interaction phase, we obtain a set of deduplicated candidate components that serve as the input for this two-step process.
-For each candidate, we issue an initial prompt to ChatGPT to generate a functional-oriented interpretation that supports both capability analysis and component determination.
-For example, given the package name com.google.mlkit.vision.objects, ChatGPT may provide the analysis result of “Google’s ML Kit API for object detection and tracking in images and videos.” In cases where components are obfuscated or highly customized, i.e., nlp.WordPieceModelPB, ChatGPT can still offer meaningful functional descriptions, e.g., “A library for tokenizing text using the WordPiece algorithm, implemented in Protocol Buffers format” which would be difficult to identify using a rule-based approach.
+After passing through the knowledge base interaction phase, we obtain a set of deduplicated candidate components that are used for further analysis.
+For each component, we issue an initial prompt to ChatGPT to produce a functional-oriented interpretation.
+For example, given the package name com.google.mlkit.vision.objects, ChatGPT may provide the analysis result of “Google’s ML Kit API for object detection and tracking in images and videos.” In cases where components are obfuscated or highly customized, i.e., nlp.WordPieceModelPB, ChatGPT can still offer meaningful functional descriptions, e.g., “A library for tokenizing text using the WordPiece algorithm, implemented in Protocol Buffers format” which would be difficult to identify using a rule-based method approach.
 Based on the generated analysis, we further prompt ChatGPT to determine whether the component falls within the scope of AI and justify its decision with a detailed explanation.
@@
 3.
-2.4. Task 4: AI Service Summarization
-Figure 3. Prompt design and examples for LLMAID
-Given the determined AI components together with their corresponding functional descriptions, we construct and employ a carefully designed ChatGPT prompt to generate a concise AI service report.
-This report synthesizes the functional interpretations of the identified components into a coherent summary of their AI capabilities, thereby providing an interpretable overview that bridges low-level implementation details with higher-level functional insights.
-Nevertheless, due to the inherent context-length limitations of ChatGPT, large language models (LLMs) in general, it is often infeasible to incorporate all component names and their corresponding analyses into a single prompt, which necessitates additional strategies to manage input complexity.
-To mitigate this constraint, we exclude low-level API components whose functionalities can generally be inferred from their package context, and instead concentrate on higher-level components.
-The remaining AI-relevant components and their interpreted AI functional descriptions are concatenated and incorporated into a carefully constructed prompt for ChatGPT, which is then used to generate the AI service report, as illustrated in Task 4 of Fig.
+2.4 Task 4: AI Service Summarization
+Given the determined AI components and their corresponding functional descriptions, we generate a concise AI service report by employing a a well-crafted ChatGPT prompt to summarize the AI capabilities of these components.
+However, due to the inherent context length limitations of ChatGPT and LLMs in general, incorporating all component names and their analyses into a single prompt is often infeasible.
+To mitigate this constraint, we omit low-level API components, i.e., their functionalities are typically inferred from their package context, and focus on higher-level components instead.
+By concatenating the remaining AI-relevant components, we prompt ChatGPT to generate the AI service report for app users, which is shown in Task 4 of Fig.
 3.
-It is worth noting that the generated report can be tailored to different audiences, such as end users, regulators, or developers, by making minor adjustments to the prompting strategy, thereby enabling context-specific interpretations and supporting diverse analytical needs.
-2.5. Experiment Configuration
-In this study, to instruct ChatGPT to execute each sub-step precisely, we carefully design natural language prompts, which are presented in Fig.
+Notably, the generated report can be adapted for different audiences, such as end users, regulators, or developers, by slightly adjusting the prompt, thereby enabling context-specific interpretations and supporting diverse analysis needs.
+2.5 Experiment Configuration
+To instruct ChatGPT to execute each sub-step precisely, we carefully design natural language prompts, which are presented in Fig.
 3.
-To further guide ChatGPT in understanding the intended analysis tasks, we provide five input–output examples to enable few-shot in-context learning(brown2020language; wang2020generalizing).
+To make ChatGPT fully understand these analysis steps, we additionally provide five input-output examples for the few-shot in-context learning brown2020language; wang2020generalizing.
 In our implementation of LLMAID, we utilize the default 4K-token context window provided by the ChatGPT API.
-However, due to ChatGPT’s rate limit, processing one app with one component per request is both time-consuming and inefficient.
-Conversely, submitting too many components in a single request risks output truncation or omission, which may cause misalignment between inputs and outputs.
+Given ChatGPT’s rate-limiting, processing one app with one component per request becomes time-consuming and inefficient.
+Moreover, submitting too many components in a single request may lead to output truncation or omission, resulting in misalignment between input and output.
 To balance task accuracy and processing efficiency, we adopt a batch size of three components per request and instruct ChatGPT to return structured outputs in JSON format.
-The model parameters are left at default values, except for the temperature and TopP settings, which are adjusted from 1.0 to 0.2 and 0.95, respectively.
-These adjustments reduce output variability while retaining a certain degree of creativity, thereby improving the consistency and reliability of the component analyses and the resulting AI service reports.
-3. Evaluation
-In this section, we evaluate the effectiveness of our proposed approach, LLMAID, by addressing the following research questions:
+Model parameters are left at default values, except for the temperature and TopP settings, which are adjusted from 1.0 to 0.2 and 0.95, respectively.
+These adjustments reduce output variability while retaining a certain degree of creativity, thereby improving the consistency and reliability of the component analyses and summary reports.
+3 Evaluation
+In this section, we evaluate the effectiveness of our approach, LLMAID, by addressing the following research questions:
   • RQ1: To what extent can LLMAID reliably and accurately determine AI capabilities in previously identified AI Apps?
     This research question aims to assess the accuracy and reliability of LLMAID in analyzing AI capabilities within previously recognized AI Android apps.
-Addressing this question validates the fundamental effectiveness of LLMAID and provides the foundation for subsequent evaluations.
+Addressing this question validates the core effectiveness of LLMAID and establishes a foundation for further evaluation.
   • RQ2: How does LLMAID compare to the state-of-the-art rule-based approach in reliably identifying AI capabilities within in-wild Android Apps?
@@
   • RQ3: How is the usefulness of the AI service summary generated by LLMAID?
-    We explore the clarity and informativeness of the AI service summaries generated by our framework.
-These summaries are intended to provide end users, developers, and regulators with an interpretable overview of the AI functionalities embedded within a given Android application.
+    We explore the clarity and informativeness of the AI service summary provided by our framework.
+This summary provides end users, developers, and regulators with an insightful overview of the AI functionalities of a given Android app.
   • RQ4: To what extent can LLMAID uncover prevalent AI capabilities and usage patterns in real-world Android Apps?
@@
 The findings offer a deeper understanding of the AI functionalities embedded in these apps and provide valuable insights to guide AI researchers and developers in designing and improving their app features.
-3.1. Dataset
+3.1 Dataset
 We perform our experiments on two separate datasets.
-The first comes from the existing literature, which was collected from AndroZoo (allix2016androzoo) and contains 56,682 AI Apps determined by a rule-based approach.
-The other is self-curated, comprising 15,517 Android apps, and reflects an open-world setting to address the absence of newer released applications, particularly those potentially integrating with emerging LLMs such as ChatGPT.
-Detailed descriptions of these two datasets are provided in the following paragraphs.
-3.1.1. Dataset 1
+The first one comes from the existing literature, which was collected from AndroZoo and contains 56,682 rule-based approach determined AI Android apps allix2016androzoo.
+The other one is self-curated, comprising 15,517 Android apps, to address the absence of recently released applications, especially those integrated with emerging LLMs such as ChatGPT.
+We present these two datasets in the following paragraphs.
+3.1.1 Dataset 1
 Li et al.
-(li2022ai) proposed a SOTA rule-based approach, AI Discriminator, to identify AI-driven Android apps via keywords matching on the dissected and decompiled Android apps.
+li2022ai proposed a SOTA rule-based approach, AI Discriminator, to identify AI-driven Android apps via keywords matching on the dissected and decompiled Android apps.
 They successfully identified 56,682 AI apps from a total of 7,259,232222The latest Android version is considered while multiple versions are stored.
 apps on AndroZoo.
-In our experiments, we leverage this dataset of 56,682 AI applications derived via the AI Discriminator approach to conduct cross-validation of our method, LLMAID.
-3.1.2. Dataset 2
+In our experiments, we utilize the dataset of 56,682 Android AI apps to perform cross-validation of our approach.
+3.1.2 Dataset 2
 To curate our dataset, we leverage AndroZoo and extract the latest release of each Android app using metadata from its associated CSV file.
-For each app, we utilize the google-play-scraper npm package to verify its current availability on Google Play Store by querying with the package name (playscraper:2023), and collect its key metadata, particularly the app description and release date.
+For each app, we utilize the google-play-scraper npm package to verify its current availability on Google Play Store by querying with the package name playscraper:2023, and collect its key metadata, particularly the app description and release date.
 In total, we identified 2,298,423 actively maintained Android apps out of a total of 23,321,823 apps available in AndroZoo at the time of our study.
 To focus our evaluation on apps with likely AI functionality, we further refined the dataset by examining the app descriptions retrieved from Google Play, selecting only those that strongly indicated the presence of AI-related features.
-Dataset 2 complements Dataset 1 by capturing newer Android apps with potential AI functionalities beyond rule-based detection, enabling a more rigorous and forward-looking assessment of LLMAID in an open-world setting.
-Existing studies have often relied on manual inspection to identify AI-enabled applications; however, such an approach is inherently time-consuming and difficult to scale to the size of popular app repositories.
-In practice, publicly available Android applications typically include short descriptions that summarize their core functionalities, offering a lightweight entry point for automated analysis.
-For example, the description of ChatGO - AI Chatbot Assistant [^3] https://play.google.com/store/apps/details?id=com.chatgpt.aichat.gpt3.aichatbot mentions that ‘Discover ChatGO - AI Chatbot Assistant, the optimized [AI assistant] that transforms the way you work, create, and connect!
-Built on cutting-edge technologies … writing, designing, summarizing, or just exploring new ideas, ChatGO is your [go-to AI companion] … and highly personalized assistance’.
-This example illustrates that AI-related applications include explicit references to AI-related terms (as highlighted in red) within their descriptions.
-Motivated by this observation, we design a two-stage process for identifying AI applications.
-In the first stage, a rule-based filter is applied to detect potential AI apps based on the presence of AI-related keywords in their descriptions.
-In the second stage, we employ ChatGPT to perform a deeper semantic analysis of the filtered descriptions, thereby verifying whether the applications are likely to incorporate AI technologies beyond superficial references to AI-related keywords.
-Specifically, the first filtering stage applies rule-based keyword extraction to identify potential AI applications based on their descriptions.
-The keyword set covers a broad spectrum, ranging from high-level terms (e.g., “AI” and “Deep Learning”) to more task-specific ones (e.g., “text classification” and “facial recognition”).
-By performing exact matches against this vocabulary, the filter effectively narrows the candidate pool, enabling subsequent analysis to concentrate on apps with higher potential relevance.
-However, the mere occurrence of such keywords does not necessarily imply that AI constitutes a core functionality of the app.
-To address this limitation, we introduce a second filtering stage powered by ChatGPT, which interprets the semantic context of app descriptions to assess whether AI techniques are likely to be employed.
-This stage improves the precision of the identification process by moving beyond simple keyword occurrence matching.
-Through this two-stage design, our approach mitigates the limitations of purely rule-based methods and more accurately detects AI-enabled Android applications.
-Nonetheless, due to semantic inflation and the frequent misuse of AI-related terms (e.g., “deep learning”), even when combining keyword matching with semantic analysis of descriptions, reliance on literal text alone remains inadequate and can lead to inaccurate identification.
-Therefore, this dataset is treated as an open-world setting and is used to evaluate the proposed approach, where incorporating implementation details enables more reliable detection of true AI apps.
-Figure 4. The distribution of collected apps over the years.
-In total, we successfully identified 15,517 Android applications (considering only the latest released version) from a corpus of 23,321,823 apps available on AndroZoo.
-Furthermore, we extracted release dates from the metadata provided on the Google Play Store and analyzed the temporal distribution of these apps, as shown in Fig.
-4.
-The results indicate that the number of released apps has grown in parallel with the increasing prevalence of AI technologies over the past decade.
+Previous studies often rely on manual inspection to identify AI apps, however, this approach tends to be time-consuming and challenging to scale.
+A publicly available Android app conventionally includes a brief description that clarifies its core functionalities.
+For example, the description of Wise AI Chatbot [^3] https://play.google.com/store/apps/details?id=com.jgapps.chatgpt mentions that ‘Wise Chat GPT Pro is the first app to utilize … ChatGPT 3.5 Turbo to provide instant and [intelligent] responses … with Wise Chat GPT Pro, our [AI] [chatbot] powered by ChatGPT 3.5 Turbo technology …’.
+This example illustrates how AI-related apps often explicitly reference AI-related terms (i.e., highlighted in red) in their descriptions.
+Motivated by this observation, we transform this step into a two-staged process.
+We begin by applying a rule-based filter to identify potential AI apps through the presence of AI-related keywords in their descriptions.
+Subsequently, we employ ChatGPT to perform a deeper semantic analysis of these filtered descriptions, further verifying whether the apps likely incorporate AI technologies.
+To be specific, the filtering stage employs rule-based keyword extraction to identify potential AI apps based on their application descriptions.
+We define a wide range of keywords from high-level functionalities (e.g., “AI” and “Deep Learning”) to low-level ones (e.g., “text classification” and “facial recognition”).
+By matching these terms exactly, the filter effectively reduces the candidate pool, allowing subsequent analysis to focus on high-potential AI apps.
+However, the mere presence of such keywords does not always indicate that AI constitutes a core functionality of the app.
+To address this limitation, we introduce a second filtering stage using ChatGPT, which evaluates whether the app is likely to employ AI techniques based on its description.
+This stage further enhances the precision of our identification process by interpreting the context of the apps’ descriptions.
+By employing the two-stage process, our approach overcomes the limitation of keyword matching methods, hence accurately detecting AI Android apps.
+Figure 4: The distribution of collected apps over the years.
+(a) Considered.
+(b) Determined.
+Figure 5: The determination of AI candidates.
+In total, we successfully identified 15,517 Android apps (only the latest released version is considered) from a total of 23,321,823 apps on AndroZoo.
+Additionally, we extracted the release date from the metadata provided on the Google Play Store and presented the distribution of these apps over the last decade in Fig.
+5.
+Our data shows that the growth in released apps aligns closely with the increasing prevalence of AI technologies during this period.
 Notably, Fig.
-4 presents only 15,055 applications, as release date metadata was unavailable for 462 apps, accounting for 2.98% (462 out of 15,517) of the total.
-This minor omission does not affect the overall observation, and the upward trend of AI technology adoption in mobile applications over the past decade remains evident.
-3.2. RQ1: To what extent can LLMAID reliably and accurately determine AI capabilities in previously identified AI Apps?
+5 includes only 15,055 apps, as release date metadata was unavailable for 462 apps, representing just 2.98% (462 out of 15,517) of the total.
+Despite this, the upward trend of employing AI technologies in mobile apps over the past ten years remains valid.
+3.2 RQ1: To what extent can LLMAID reliably and accurately determine AI capabilities in previously identified AI Apps?
 In this research question, we examine whether LLMAID can effectively identify AI capabilities within a set of pre-labeled AI applications.
-For this purpose, we employ Dataset 1 as our evaluation benchmark.
-Since Dataset 1 identifies AI apps but does not provide ground-truth labels for their underlying AI components, we randomly sample a subset of 382 applications and conduct human validation to assess the accuracy of LLMAID.
+We employ Dataset 1 as our evaluation benchmark.
+However, since Dataset 1 only determines AI apps but does not provide ground-truth labels for their detailed underlying AI components, we randomly sample a subset of 382 AI apps and perform human validation to assess LLMAID.
 The sample size is calculated based on the popular online Sample Size Calculator  [^4] https://www.surveysystem.com/sscalc.htm with a confidence level of 95% and a margin of error of 5%.
-We first present the quantity distribution of three representative types of components, including packages, APIs, and HTTPS requests in Fig.
+We first present the quantity distribution of three representative types of components, packages, APIs, and HTTPS requests in Fig.
 5.
@@
 The whitelist of non-AI components comprises the official Java and Android kernel-related packages and API signatures.
-After performing AI capability analysis and determination, the median number of AI-related packages, APIs, and HTTPS requests drops to 9.5, 33, and 0, respectively.
+After performing AI capability analysis, the median number of AI-related packages, APIs, and HTTPS requests drops to 9.5, 33, and 0, respectively.
 Notably, packages are filtered more aggressively than APIs, which is expected given that many AI-specific APIs are concentrated within a smaller set of AI-related packages.
 In contrast, HTTPS requests are sparsely detected, primarily due to the relatively recent adoption of AI-related network calls and their limited presence in the collected dataset.
-(a) Considered.
-(b) Determined.
-Figure 5. The determination of AI candidates.
 To evaluate the reliability of LLMAID in identifying AI components within the sampled dataset, we conducted a manual annotation study involving two independent authors, who independently labeled a total of 49,371 components extracted from the 382 sampled AI apps.
-When disagreements arise, dedicated meetings with an additional author are held to facilitate consensus.
+When disagreements arise, dedicated meetings involving an additional author are held to facilitate consensus.
 Each component was classified as either an AI component or a non-AI component.
@@
 Following human validation, LLMAID achieved a precision of 98.05% and a recall of 93.31%, indicating strong performance in correctly detecting AI components.
-Furthermore, the inter-annotator agreement, measured using Cohen’s unweighted kappa coefficient (cohenkappa), was 0.82, indicating substantial consistency among the human evaluators.
+Furthermore, inter-annotator agreement measured using Cohen’s unweighted kappa coefficient cohenkappa was 0.82, reflecting substantial consistency among human evaluators.
 These results collectively demonstrate the high accuracy and robustness of LLMAID in detecting AI components.
 In addition, we validated the reliability of LLMAID in identifying AI apps.
-We defined an app as an AI app if it contains one or more AI components determined by LLMAID.
-Our results showed that LLMAID correctly identified 378 out of 382 sampled applications as AI apps, corresponding to an accuracy of 99.0% (378/382).
-For the remaining four applications, we further manually inspected them and observed that the only potential AI indicators were filenames with the .model suffix.
+We defined an app is considered an AI app if it contains one or more AI components determined by LLMAID.
+Our results showed that LLMAID can detect 378 out of 382 sample apps as AI apps, indicating an accuracy of 99.0% (372/382) is achieved.
+We further analyzed the remaining four apps and observed that the only potential AI indicators were filenames with the .model suffix.
 Without further complementary information, such as associated APIs, packages, or functional descriptions, LLMAID was unable to classify these apps as AI apps, hence excluding them from the set of AI apps.
-Table 1. Comparison of the reliability of LLMAID between the implementations of analysis-then-detection and detection-then-analysis.
-Analysis-then-Detection Detection-then-Analysis
+Table 1: Comparison of the reliability of LLMAID between the implementations of analysis-then-detection and detection-then-analysis.
+Analysis/Detection Detection/Analysis
 AI Candidate
@@
     Average running time for an Android APK in seconds.
-Although the design of Task 3.1 (AI Capability Analysis) and Task 3.2 (AI Component Determination) is inspired by Chain-of-Thought prompt (wei2022chain), this sequence may not be optimal for LLMAID.
-A more efficient alternative is to first determine whether a component is AI-related or not and then analyze and interpret only those classified as AI components, effectively reversing the order of Tasks 3.1 and 3.2.
-Under the same evaluation setup on the 382 sampled apps, we find that this detection-then-analysis approach reduces the average processing time per app by 65%, as summarized in Table 1.
-However, the number of AI components identified using this approach is notably lower than that obtained with the original sequential implementation (Analysis-then-Detection).
+Although the design of Task 3.1 (AI Capability Analysis) and Task 3.2 (AI Component Determination) is inspired by Chain-of-Thought prompt wei2022chain, it may not be optimal for LLMAID.
+A more efficient alternative is to first classify the components and then analyze only those identified as AI components, reversing the order of Tasks 3.1 and 3.2.
+Following the same evaluation setup on the 382 sampled apps, we find that this detection-then-analysis approach reduces the processing time per app by 65%, as shown in Table 1.
+The number of identified AI components is significantly lower than that of the original sequential implementation.
 Based on the overall results, we suggest that detection-then-analysis can only be used when developers are interested in finding AI apps.
-If developers seek to understand all possible AI capabilities of an app, we foresee that the original LLMAID with analysis-then-detection is more effective in capturing AI candidates.
+If developers seek to understand all possible AI capabilities of an app, we foresee that the original LLMAID with analysis-then-analysis is more effective in capturing AI candidates.
 RQ1 Findings Human validation suggests that LLMAID achieves high precision (98.05%) and recall (93.31%) in detecting AI components and AI apps.
 Although reversing the sequence of Task 3.1 and Task 3.2 could improve the efficiency of LLMAID, it compromises reliability compared to the original sequential implementation.
-Table 2. Number of AI and Non-AI apps classified by AI Discriminator and LLMAID.
+3.3 RQ2: How does LLMAID compare to a state-of-the-art rule-based approach in reliably identifying AI capabilities within in-wild Android Apps?
+In this research question, we evaluate how reliable LLMAID is in determining AI capabilities and compare it to the SOTA approach, AI Discriminator li2022ai; aiapp:2023.
+To efficiently perform the evaluation, we select a representative subset of Dataset 2, focusing on the apps published in the recent five years, and randomly sampling 30% apps in each year to preserve the same quantity distribution.
+Finally, we ended up with 3,819 $(1980\times 30\%+2336\times 30\%+2326\times 30\%+3053\times 30\%+3044\times 30\%)$ apps for the experiment.
+Following the same experimental setup as in RQ1, we compute the precision and recall for AI component identification, and LLMAID achieves a precision of 97.32% and a recall of 91.01% in identifying AI components.
+The inter-annotator reliability, measured by Cohen’s unweighted kappa coefficient, is 0.80, indicating a high level of agreement with the results generated by LLMAID.
+Table 2: Number of AI and Non-AI apps classified by AI Discriminator and LLMAID.
 AI Discriminator LLMAID
-AI Non-AI AI Non-AI
-555 3264 1366 2453
-Table 3. Detected AI Apps by AI Discriminator and LLMAID
+AI Apps Non-AI Apps AI Apps Non-AI Apps
+603 3216 1366 2453
+Table 3: Detected AI Apps by AI Discriminator and LLMAID
 AI Discriminator
@@
 9 546 820
-3.3. RQ2: How does LLMAID compare to a state-of-the-art rule-based approach in reliably identifying AI capabilities within in-wild Android Apps?
-In this research question, we evaluate how reliable LLMAID is in determining AI capabilities and compare it to the SOTA approach, AI Discriminator (li2022ai; aiapp:2023).
-To efficiently perform the evaluation, we select a representative subset of Dataset 2, focusing on the apps published in the recent five years, and randomly sampling 30% apps in each year to preserve the same quantity distribution.
-Finally, we ended up with 3,819 $(1980\times 30\%+2336\times 30\%+2326\times 30\%+3053\times 30\%+3044\times 30\%)$ apps for the experiment.
-Following the same experimental setup as in RQ1, we compute the precision and recall for AI component identification.
-LLMAID achieves a precision of 97.32% and a recall of 91.01%, demonstrating strong performance in identifying AI components.
-The inter-annotator reliability, measured using Cohen’s unweighted kappa coefficient, is 0.80, which reflects a high level of agreement with the results generated by LLMAID.
-Table 3.2 presents the experimental results for AI App identification, comparing the performance of AI Discriminator with that of LLMAID.
-Among the total of 3,819 experimental Android Apps, 555 of them are classified by AI Discriminator as AI Apps, while 1,366 are recognized as AI Apps by LLMAID, which is 242% more than AI Discriminator.
+Table 3.3 presents the experimental results of AI App identification, comparing the performance of AI Discriminator and LLMAID.
+Among the total of 3,819 experimental Android Apps, 603 of them are classified by AI Discriminator as AI Apps, while 1,366 are recognized as AI Apps by LLMAID, which is 242% more than AI Discriminator.
 Table 3.3 shows the number of AI Apps identified by AI Discriminator and LLMAID.
 It clearly demonstrates that 546 Android Apps, accounting for 90.55% $(546/603)$ of those identified by AI Discriminator and 39.97% $(546/1366)$ of those confirmed by LLMAID, are commonly recognized as AI Apps by both approaches.
-To manually inspect the nine AI apps identified solely by AI Discriminator, multiple authors downloaded the APK files from AndroZoo and dissected them with Jadx (jadx2023).
+To manually inspect the nine AI apps identified solely by AI Discriminator, multiple authors downloaded the APK files from AndroZoo and dissected them with Jadx jadx2023.
 Two apps are recognized as AI apps because only one special tflite file is being detected, while the remaining seven apps are classified as AI apps because of keywords, such as “DecisionTreeClassifier” and “KNeighborsClassifier” in their descriptions of json files.
-However, the description contents in the JSON files alone, without supporting code-level detection, may not imply the actual implementation of AI capabilities and therefore should not be regarded as convincing AI evidence.
-To verify the 820 apps only classified by LLMAID, we randomly sampled 86 apps (with a confidence level of 95% and a margin of error of 10%) and cross-validated them using the decompiler tool, Jadx (jadx2023), with multiple authors involved.
-In the course of this cross-validation, we observed that OpenAI APIs were sometimes directly invoked within Java code, whereas in other cases developers implemented these API calls through native libraries (i.e.,.so files).
-Such implementations of cloud-based AI capabilities could not be detected by AI Discriminator due to the limitations of its rule set.
-In the end, we confirmed that 83 of the total of 86 sampled apps are classified by LLMAID as truly AI apps, showcasing a high level of identification precision.
+However, the description contents in the JSON files may not imply the implementation of AI capability and therefore should not be counted as convincing AI evidence.
+To verify the 820 apps only classified by LLMAID, we randomly sample 86 apps (with a confidence level of 95% and a margin of error of 10%) and cross-validated by multiple authors with the decompiler tool, Jadx jadx2023.
+During cross-validation, we observe that OpenAI APIs are sometimes directly invoked within Java code, whereas other developers implement these API invocations via native libraries (i.e.,.so files).
+This API invocation implementation of the cloud AI capabilities could not be detected by the AI discriminator, given its limited rule set.
+In the end, we confirm that 83 of the total of 86 sampled apps are classified by LLMAID as AI apps, showcasing a high level of identification precision.
 RQ2 Findings High precision (97.32%) and recall (91.01%), and the high inter-annotator reliability about the identification of AI components reconfirm the effectiveness of LLMAID for in-wild apps.
 Furthermore, the experimental result of detecting more than twice as many AI apps as the rule-based approach instantiates the outperformance of LLMAID against the SOTA one.
-Figure 6. Description of Copy.ai (the left) and its AI service summarization (the right) generated by our framework.
-3.4. RQ3: How is the usefulness of the AI service summary generated by LLMAID?
-In this research question, we seek to evaluate the usefulness of the AI summary, which is generated in Task 4 of LLMAID based on the app’s implementation details.
-Specifically, we compare the generated AI summaries with the developer-provided descriptions on the app store, using this comparison as a diagnostic tool to reveal the shortcomings of the developer-provided information and to enable a more accurate understanding and evaluation of AI functionalities.
+3.4 RQ3: How is the usefulness of the AI service summary generated by LLMAID?
+Figure 6: Description of Copy.ai (the left) and its AI service summarization (the right) generated by our framework.
+In this research question, we seek to evaluate the usefulness of the AI summary, which is generated in Task 4 of LLMAID.
 We conducted a user study by recruiting six developers, each with at least three years of experience in Android app analysis.
 Specifically, we randomly selected 20 Android apps identified as AI apps in RQ2, and then generated their AI summaries.
-For each app, we paired each LLMAID generated AI summary with its corresponding app description retrieved from the Google Play store via the google-play-scraper (playscraper:2023), and displayed them side-by-side to participants, as shown in Fig.
+For each app, we paired each LLMAID generated AI summary with its corresponding app description retrieved from the Google Play store via the google-play-scraper playscraper:2023, and displayed them side-by-side to participants, as shown in Fig.
 6.
@@
 Each dimension was assessed using a 10-point Likert scale, with scores ranging from 1 (indicating poor quality) to 10 (indicating high quality).
-Table 4. Average and median scale values voted for sampled Android apps by six participants. “Recom.” stands for Recommendation, “devs” stands for developers, “Med.” stands for median.
+Table 4: Average and median scale values voted for sampled Android apps by six participants. “Recom.” stands for Recommendation, “devs” stands for developers, “Med.” stands for median.
 Summary Type Accuracy Completeness Consistency Recom. app users Recom. app devs.
@@
 RQ3 Findings The user study shows that the AI service summary generated by LLMAID is constantly preferred by practitioners in terms of accuracy, completeness, consistency, and recommendation.
-Table 5. Different types of identified AI components.
-Data Type
-Package API
+3.5 RQ4: To what extent can LLMAID uncover prevalent AI capabilities and usage patterns in real-world Android Apps?
+Table 5: Different types of identified AI components.
+Data
+Type
+Package
+API
 Model
-HTTP Requests
+HTTP
+Requests
 Others Total
 #. items 10,285 644 120 32 11,081
-3.5. RQ4: To what extent can LLMAID uncover prevalent AI capabilities and usage patterns in real-world Android Apps?
+Figure 7: Detailed tasks and corresponding AI domains of the AI components.
 Regarding this research question, we seek to classify the specific domains and tasks of the identified AI components in the determined AI apps in RQ2.
@@
 The second largest category of AI components is model, which refers to on-device AI model files, such as .tflite and .caffemodel.
-Apart from package names, APIs, and on-device AI models, we observe that some apps rely on HTTP requests (e.g., OpenAI APIs) to access cloud-based AI services, which has rarely been discussed in previous studies (xu2019first; li2022ai).
+Apart from package names, APIs, and on-device AI models, we observe that some apps rely on HTTP requests (e.g., OpenAI APIs) to access cloud-based AI services, which has rarely been discussed in previous studies xu2019first; li2022ai.
 The remaining components exhibiting limited shared patterns are collectively classified under the category of Others.
-Figure 7. The distribution of tasks (inner ring) and their corresponding AI domains (outer ring) for the AI components.
+Table 6: Number of Android apps in different domains. DADA: Data Analysis; CVCV: Computer Vision; NLPNLP: Natural Language Processing; ASPASP: Audio and Speech Processing; ARAR: Augmented Reality.
+Domains DA CV NLP ASP AR Others
+#. Apps 643 634 43 30 2 14
 With these identified AI components, we first reviewed the analysis results to design a classification scheme that captures a wide range of AI domains and tasks.
@@
 The area of each colored sector represents its corresponding proportion.
-At the top level, AI components are classified into six common AI domains, including Computer Vision (54.80%), Data Analysis (26.85%)  [^5] Data analysis and processing denote statistical operations within AI frameworks., Natural Language Processing (13.38%), Others (2.75%), Audio and Speech Processing (1.90%), and Augmented Reality (0.32%).
+At the top level, AI components are classified into six common AI domains, including Computer Vision (54.80%), Data Analysis (26.85%), Natural Language Processing (13.38%), Others (2.75%), Audio and Speech Processing (1.90%), and Augmented Reality (0.32%).
 The category Others contains AI components that are difficult to classify clearly.
@@
 The second largest AI domain is Data Analysis, which is occupied mostly by the task of Data Processing (86.69%), such as tensor operations and input/output data parsing.
-Table 6. Number of Android apps in different domains.
-Domains
-Data Analysis
-Computer Vision
-Natural Language
-Processing (NLP)
-Audio & Speech
-Processing (ASP)
-Augmented Reality
-Others
-#. Apps 643 634 43 30 2 14
 Based on classified AI components, we further categorize the domains and tasks at the app level.
@@
 In contrast, the augmented reality oriented ones are the least provided due to the limited number of compelling real-world application scenarios, despite the concept’s novelty and growing popularity.
-4. Discussion
+4 Discussion
 We now discuss the inherent limitations (Section  4.1) and possible implications for both researchers and practitioners based on the outcomes of our study (Section 4.2).
-4.1. Limitations
-4.1.1. LLM Hallucination
+4.1 Limitations
+4.1.1 LLM Hallucination
 Hallucination, referring to incorrect or fabricated responses to user inquiries, and inconsistency in responding to identical queries are inherent limitations of LLMs.
@@
 Our manual inspection of the sampled experimental results indicates that LLMAID delivers reliable and accurate analyses, demonstrating its robustness despite the inherent limitations of LLMs.
-Figure 8. The percentage of obfuscated Java file names and their directory names.
-4.1.2. Obfuscations in Android apps
-Obfuscation techniques, such as string encryption (e.g., the obfuscated API: <<d.f.e.x.b: void <<init>> (int)>>) and Java reflection, are widely adopted in Android apps to safeguard intellectual property against reverse engineering, as extensively evidenced by existing research (dong2018understanding; glanz2020hidden).
-We have attempted to replicate some deobfuscation artifacts, such as StringHound (glanz2020hidden) and simplify (guillot2010automatic), to recover the obfuscated items in the Android apps in our study.
+Figure 8: The percentage of obfuscated Java file names and their directory names.
+4.1.2 Obfuscations in Android apps
+Obfuscation techniques, such as string encryption (e.g., the obfuscated API: <<d.f.e.x.b: void <<init>> (int)>>) and Java reflection, are widely adopted in Android apps to safeguard intellectual property against reverse engineering, as extensively evidenced by existing research dong2018understanding; glanz2020hidden.
+We have attempted to replicate some deobfuscation artifacts, such as StringHound glanz2020hidden and simplify guillot2010automatic, to recover the obfuscated items in the Android apps in our study.
 However, these tools either failed to execute successfully in our environment or produced unsatisfactory results.
@@
 While obfuscation may inevitably introduce some limitations, the substantial portion of unobfuscated implementation details still enables us to extract meaningful insights and supports the reliability of our overall analysis and summarization.
-4.2. Implications
-4.2.1. On-Device and cloud AI integration in Android apps.
-The growing integration of on-device models with cloud-based LLM APIs in Android applications introduces new opportunities and challenges that warrant systematic investigation.
-While such hybrid architectures promise enhanced proactivity, context-awareness, and personalization, they also raise unresolved issues of reliability, privacy, and ethical responsibility.
-Future research should extend LLMAID to analyze how Android apps orchestrate local and cloud intelligence, assess resilience under failures or connectivity disruptions, and examine the privacy–utility trade-offs inherent in offloading user data to remote services.
-Equally important is understanding the evolution of developer practices in adopting hybrid AI features, from SDK/API usage to prompt design and deployment strategies, as well as addressing transparency and fairness concerns in user-facing AI behaviors.
-Advancing these directions will be essential for establishing trustworthy, adaptive, and ethically responsible AI integration in the Android ecosystem.
-4.2.2. Security mitigation for software artifacts.
+4.2 Implications
+4.2.1 Security mitigation for software artifacts.
 Our study reveals that by asking LLM about the implementation basics, it could respond with accurate, at least to some extent, descriptions of whether they are AI-related or not.
@@
 Consequently, more reliable and secure artifacts could be achieved.
-4.2.3. Software artifacts analysis for researchers.
+4.2.2 Software artifacts analysis for researchers.
 Our research on the summarization of software functionality in RQ3 validates the potential of LLM in code interpretation and summarization, specifically within the context of Java as a single programming language.
@@
 Detailed analysis of interactions between different languages and careful design of prompts are necessary to make full use of the powerful potential offered by LLMs, which also paves a promising research direction for our research communities.
-4.2.4. Advanced AI identification approaches & questionnaire with app developers.
+4.2.3 Advanced AI identification approaches & questionnaire with app developers.
 In the AI app identification process, we found that some self-claimed AI Android apps, despite describing advanced AI features in their Play Store listings, contained no detectable AI components in the released APKs.
-A possible reason is the use of Android App Bundles (appbundle:2023), which deliver only essential components at the installation time and provide the necessary AI functions in the on-demand parts, preventing full inspection of the implementation details.
+A possible reason is the use of Android App Bundles appbundle:2023, which deliver only essential components at the installation time and provide the necessary AI functions in the on-demand parts, preventing full inspection of the implementation details.
 To enable comprehensive analysis, it is imperative to adopt a new approach capable of harvesting all parts of an Android app, not only the basic ones but also those loaded on demand.
@@
 Therefore, a questionnaire with these popular Android app developers would provide us with a few great insights for further app analysis.
-5. Related Work
-5.1. AI Mobile Apps Study
+5 Related Work
+5.1 AI Mobile Apps Study
 Xu et al.
-(xu2019first) first proposed an APK analysis tool, DL Sniffer to empirically investigate the use of deep learning techniques in Android apps.
+xu2019first first proposed an APK analysis tool, DL Sniffer to empirically investigate the use of deep learning techniques in Android apps.
 The DL Sniffer dissects the released APK and identifies the utilization of 16 popular deep learning frameworks via two separate modules.
-The first one recognizes specific strings (e.g., “TF_AllocateTensor”) in the section rodata of native shared libraries (i.e., file ends with “.so”) via some powerful Linux command (readelf:2023; objdump:2023).
-The other one identifies AI special APIs in DEX binary files by de-compiling the binary file into smali code via (dex2jar:2023).
 Experiments on the large-scale dataset containing 16,500 Android apps reveal that top apps are AI techniques early adopters, especially photo beauty apps, and the on-device AI is actually lighter than practitioners expect.
+In addition, the DL framework developers should take more time to optimize and protect frameworks.
 Sun et al.
-(sun2021mind) proposed ModelXRay to determine whether on-device AI frameworks are adopted and to check if the adopted framework is protected or encrypted in a large dataset containing 46,753 Android apps.
+sun2021mind proposed ModelXRay to determine whether on-device AI frameworks are adopted and to check if the adopted framework is protected or encrypted in a large dataset containing 46,753 Android apps.
 Their experimental results show that 41% Android apps do not protect the adopted DL models, and even some of the protected models still have a weak status, which would result in the leakage of private information and economic losses.
 Li et al.
-(li2022ai) conducted an exploratory study and released a tool, AI Discriminator, to analyze the adoption of AI technologies in Android apps.
-They first refined the rules used in the identification of AI app and successfully recognized 56,682 AI apps from a total of 7,259,232 mobile apps provided on AndroZoo (allix2016androzoo).
-With the historical versions of apps considered, they concluded that the AI models utilized in Android apps are not updated regularly and are not protected carefully.
-What’s more, AI apps always request access to various types of sensitive user data.
+li2022ai conducted an exploratory study and released a tool, AI Discriminator, to analyze the adoption of AI technologies in Android apps.
+They first refined the rules used in the identification of AI app and successfully recognized 56,682 AI apps from a total of 7,259,232 mobile apps provided on AndroZoo allix2016androzoo.
+They concluded that the AI models utilized in Android apps are not updated regularly and are not protected carefully.
 In addition, they analyzed reviews for the determined AI apps and revealed that users always express positive attitudes towards the adoption of advanced AI technologies.
-5.2. LLMs in Software Engineering
+5.2 LLMs in Software Engineering
 Feng et al.
-(feng2023prompting) tried to reproduce Android bugs with the help of LLMs.
+feng2023prompting tried to reproduce Android bugs with the help of LLMs.
 They proposed a lightweight bug replay tool, AdbGPT, to automatically rerun Android APKs and trigger bugs according to collected bug reports.
-The AdbGPT first utilizes prompt engineering , working as real app developers, to extract the necessary steps for bug replay from the given bug reports.
-And then prompts LLMs to run Android apps under the guidance extracted previously.
-Their extensive experiments reveal that AdbGPT could obtain a higher accuracy in necessary step extraction compared with the state-of-the-art traditional approaches and finally successfully reproduce 81.3% bugs, which is much better than baselines.
+Their extensive experiments reveal that AdbGPT could successfully reproduce 81.3% bugs, which is much better than baselines.
 Their approach demonstrates the effectiveness of LLMs in assisting traditional software research problems.
 Tian et al.
-(tian2023chatgpt) focused on the research tasks of code generation, program repair, and code summarization and conducted an exploratory study to investigate performance in the aforementioned practices.
-They curated datasets from LeetCode 2016-2020 and LeetCode 2022, and constructed prompts to generate code with problem descriptions, repair buggy code implementations, and summarize code intentions.
+tian2023chatgpt focused on the research tasks of code generation, program repair, and code summarization and conducted an exploratory study to investigate performance in the aforementioned practices.
 The experimental results not only reveal that ChatGPT is effective in common software practices but also unveil that limitations exist in LLMs, such as long and detailed prompt descriptions that would hinder the effectiveness of ChatGPT.
 Ma et al.
-(ma2023scope) conducted a study to touch on the limitations of the interpretability of ChatGPT in three types of software engineering jobs, including understanding syntax, static behaviors, and dynamic behaviors.
+ma2023scope conducted a study to touch on the limitations of the interpretability of ChatGPT in three types of software engineering jobs, including understanding syntax, static behaviors, and dynamic behaviors.
 Performing ChatGPT on different programming languages, the authors concluded that ChatGPT does have the ability to construct AST but it still produces incorrect or fake facts with respect to dynamic semantic understanding.
-It is necessary for researchers to propose effective methods to verify the correctness of the generated outputs from ChatGPT.
 Li et al.
-(li2023hitchhiker) carried out an LLM-assisted static analysis with respect to use-before-initialization (UBI) bugs for C/C++ programs.
-They proposed an automatic agent that interacts with both a static analysis tool and an LLM.
-With the state-of-the-art static analysis tool and meticulous design of prompts, the authors could successfully overcome several challenges, such as modeling bugs specific to the problem, which are hard to manage within separate research directions.
+li2023hitchhiker carried out an LLM-assisted static analysis with respect to use-before-initialization (UBI) bugs for C/C++ programs.
 They conducted experiments on real potential UBI bugs, achieved higher precision and recall, and even identified 13 previously untouched ones, revealing the effectiveness of LLM in assisting software engineering problems.
-6. Conclusion
+6 Conclusion
 In this paper, we have proposed an automated approach LLMAID, which employs program analysis techniques to extract detailed implementation components from Android apps and utilizes a large language model (ChatGPT in this study) to interpret and summarize the underlying AI services.
@@
 Furthermore, with these findings, we outline several promising research directions to inform and inspire future work in this area.
-Data Availability: The source code and datasets are all made publicly available in our artifact package via the following link: https://zenodo.org/records/15522834
+7 Declarations
+Funding: Not applicable.
+Ethical approval: Not applicable.
+Informed consent: Not applicable.
+Author Contributions: Pei Liu led the project, oversaw its administration, and drafted the original manuscript.
+Terry Yue Zhuo developed and implemented the framework.
+Jiawei Deng curated the dataset and performed the experimental analyses.
+Zhenchang Xing, Qinghua Lu, Xiaoning Du conceived and designed the study and contributed to manuscript writing.
+Hongyu Zhang provided critical revisions and editorial improvements.
+Data Availability Statements: The source code and datasets are all made publicly available in our artifact package via the following link: https://zenodo.org/records/15958951
+Conflict of Interest: The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Clinical trial number: Not applicable.